### v1

# Exploring Consistency in Graph Representations: from Graph Kernels to Graph Neural Networks

Abstract
Graph Neural Networks (GNNs) have emerged as a dominant approach in graph representation learning, yet they often struggle to capture consistent similarity relationships among graphs. While graph kernel methods such as the Weisfeiler-Lehman subtree (WL-subtree) and Weisfeiler-Lehman optimal assignment (WLOA) kernels are effective in capturing similarity relationships, they rely heavily on predefined kernels and lack sufficient non-linearity for more complex data patterns. Our work aims to bridge the gap between neural network methods and kernel approaches by enabling GNNs to consistently capture relational structures in their learned representations. Given the analogy between the message-passing process of GNNs and WL algorithms, we thoroughly compare and analyze the properties of WL-subtree and WLOA kernels. We find that the similarities captured by WLOA at different iterations are asymptotically consistent, ensuring that similar graphs remain similar in subsequent iterations, thereby leading to superior performance over the WL-subtree kernel. Inspired by these findings, we conjecture that the consistency in the similarities of graph representations across GNN layers is crucial in capturing relational structures and enhancing graph classification performance. Thus, we propose a loss to enforce the similarity of graph representations to be consistent across different layers. Our empirical analysis verifies our conjecture and shows that our proposed consistency loss can significantly enhance graph classification performance across several GNN backbones on various datasets. footnotetext: Code: https://github.com/GraphmindDartmouth/Graph-consistency
1 Introduction
Graph classification tasks are extensively applied across multiple domains, including chemistry (Liu et al., 2022; Xu et al., 2023), bioinformatics (Yan et al., 2019; Li et al., 2023a, b), and social network analysis (Ying et al., 2018; Wang et al., 2024). Graph neural networks (GNNs) (Kipf and Welling, 2017; Xu et al., 2019; Velickovic et al., 2018; Huang et al., 2024) have emerged as the predominant approach for performing graph classification, owing to their ability to extract rich representations from various types of graph data. A typical GNN employs the message-passing mechanism (Gilmer et al., 2017), where node features are propagated and aggregated across connected nodes. This process effectively captures local tree structures, enabling the differentiation between various graphs. However, GNNs often struggle to preserve relational structures among graphs, resulting in inconsistent relative similarities across the layers. As shown in Figure 1, graphs with higher relative similarity in one layer may exhibit reduced similarity in the subsequent layer. This phenomenon arises from the limitations of cross-entropy loss, which fails to preserve relational structures, as it forces graphs within the same class into identical representations.
Figure 1: Cosine similarity of three molecules from the NCI1 dataset, evaluated using graph representations from three consecutive GIN layers. Common GNN models fail to preserve relational structures across the layers.
Graph kernel methods, on the other hand, are designed to capture similarities between graphs and utilize these similarities for classification tasks. For instance, subgraph-pattern approaches (Shervashidze et al., 2009; Costa and Grave, 2010; Kriege et al., 2020) compare graphs by counting the occurrences of fixed-size subgraph motifs. Other methods compare sequences of vertices or edges encountered during graph traversals (Borgwardt and Kriegel, 2005; Kashima et al., 2003; Zhang et al., 2018). Among all graph kernels, two notable ones are the Weisfeiler-Lehman subtree (WL-subtree) kernel (Shervashidze et al., 2011) and the Weisfeiler-Lehman optimal assignment (WLOA) kernel (Kriege et al., 2016). They are found to have comparable performance to simple GNNs (Nikolentzos et al., 2021). The WL-subtree kernel iteratively relabels graphs using the Weisfeiler-Lehman algorithm (Weisfeiler and Lehman, 1968) and constructs a kernel based on the number of occurrences of each label. The WLOA kernel uses the same relabeling scheme but computes a matching between substructures to reveal structural correspondences between the graphs.
While effective in capturing relative graph similarity, kernel methods rely on predefined kernels and exhibit insufficient non-linearities, limiting their ability to capture complex patterns in high-dimensional data. Additionally, kernel methods are computationally costly, making them unsuitable for handling large datasets and consequently limiting their overall applicability.
In this work, we aim to bridge the gap between kernel methods and GNN models. Given the iterative nature of GNNs, we study a class of kernels which are induced from graph representations obtained through an iterative process and name them iterative graph kernels (IGK). Within this framework, we define the consistency property, which ensures that similar graphs remain similar in subsequent iterations. Our analysis demonstrates that kernels with this property yield better classification performance. Furthermore, we find that this property sheds light on why the WLOA kernel outperforms the WL-subtree kernel. The WLOA kernel asymptotically demonstrates consistency as the iteration goes to infinity, whereas the WL-subtree kernel does not exhibit this behavior. Inspired by these findings and the analogy between message-passing GNNs and the WL-subtree kernel (Xu et al., 2019), we hypothesize that this principle is also applicable to GNNs. To explore this, we introduce a novel loss function designed to align the ranking of graph similarities across GNN layers. The aim is to ensure that the relational structures of the graphs are preserved and consistently reflected throughout the representation space of these layers. We validate this hypothesis by applying our loss function to different GNN backbones across various graph datasets. Extensive experiments demonstrate that our proposed model-agnostic consistency loss improves graph classification performance comprehensively.
In summary, the main contributions of this work are as follows:
  • Novel perspective: We present a novel perspective on understanding the graph classification performance of GNNs by analyzing the similarity relationships captured by different layers.
  • New insights: We are the first to introduce and formalize the consistency principle within both kernel-based and GNN methods for graph classification tasks. Additionally, we provide theoretical proofs explaining how this principle enhances the performance.
  • Simple yet effective method: Empirical results demonstrate that the proposed consistency loss universally enhances performance across a wide range of base models and datasets.
2 Preliminaries
In this section, we begin by introducing the notations and definitions used throughout the paper. Next, we provide an introduction to the fundamentals of Weisfeiler-Lehman isomorphism test, GNNs and graph kernels.
2.1 Notations and Definitions
Let $\mathcal{G}$($\mathcal{V}$, $\mathcal{E}$, $\mathbf{X}$) be an undirected and unweighted graph with $N$ nodes, where $\mathcal{V}$ denotes the node set, $\mathcal{E}$ denotes the edge set, and $\mathbf{X}$ denotes the feature matrix, where each row represents the features of a corresponding node. The neighborhood of a node $v$ is defined as the set of all nodes that connected to $v$: $\mathcal{N}(v)=\{u|(v,u)\in\mathcal{E}\}.$ In a graph dataset, each graph $\mathcal{G}_{i}$ is associated with a label $\mathcal{Y}_{i}$, which is sampled from a label set ${\mathcal{L}}$. In this paper, we focus on the graph classification task, where a model $\phi$ is trained to map each graph to its label.
2.2 Weisfeiler-Lehman Isomorphism Test
We first introduce the Weisfeiler-Lehman isomorphism test (Weisfeiler and Lehman, 1968), which can be used to distinguish different graphs and is closely related to the message-passing process of GNNs (Xu et al., 2019). The WL algorithm operates by iteratively relabeling the colors of vertices in the graph. Initially, all vertices are assigned the same color $C_{0}$. In iteration $i$, the color of a vertex $v$ is updated based on its current color $C_{v,i-1}$ and the colors of its neighbors $\left\{C_{u\in\mathcal{N}({v}),i-1}\right\}$. The update is given as follows:
$C_{{v},i}=f_{c}^{i}\left(\left\{C_{{v},i-1},\left\{C_{u\in\mathcal{N}({v}),i-1}\right\}\right\}\right)$
where $f_{c}^{i}$ is an injective coloring function that maps the multisets to different colors at iteration $i$.This process continues for a predefined number of iterations or until the coloring stabilizes (i.e., the colors no longer change).
2.3 Graph Neural Network
Most GNNs adopt the message-passing framework (Gilmer et al., 2017), which can be viewed as a derivative of the Weisfeiler-Lehman coloring mechanism. Specifically, let $\bm{h}_{v}^{(k-1)}$ represent the feature vector of node $v$ at the $(k-1)$-th iteration. A GNN computes the new feature for $v$ by aggregating the representations of itself and its neighboring nodes $u\in\mathcal{N}(v)$ as follows:
$\bm{h}_{v}^{(k)}=\texttt{UPDATE}^{(k)}\left(\bm{h}_{v}^{(k-1)},\bm{m}_{v}^{(k)}\right)\text{, where }\bm{m}_{v}^{(k)}=\texttt{AGGR}^{(k)}\left(\left\{\bm{h}_{u}^{(k-1)}:u\in\mathcal{N}(v)\right\}\right)$
The initial node representations $\bm{h}_{v}^{(0)}$ are set to the raw node features $\mathbf{X}_{v}$. At the $k$-th iteration, the aggregation function $\texttt{AGGR}^{(k)}(\cdot)$ computes the messages $\bm{m}_{v}^{(k)}$ received from neighboring nodes. Subsequently, the update function $\texttt{UPDATE}^{(k)}(\cdot)$ computes a new representation for each node by integrating the neighborhood messages $\bm{m}_{v}^{(k)}$ with its previous embedding $\bm{h}_{v}^{(k-1)}$. After $T$ iterations, the final node representations are combined into a graph representation using a readout function:
$\bm{h}_{G}=\texttt{READOUT}\left(\left\{\bm{h}_{v}^{(T)}\mid v\in\mathcal{V}\right\}\right).$
The readout function, essentially a set function learned by the neural network, commonly employs AVERAGE or MAXPOOL.
2.4 Graph Kernel
[…]
Graph kernel methods apply the kernel approaches to graph data, typically defined using the $\mathcal{R}$-convolution framework (Haussler et al., 1999; Bause and Kriege, 2022). Consider two graphs, $\mathcal{G}$ and $\mathcal{G}^{\prime}$. The key idea is to decompose the graphs into substructures using a predefined feature map function $\phi$, and then compute the kernel value by taking the inner product in the feature space: $\mathbb{K}(\mathcal{G},\mathcal{G}^{\prime})=\langle\phi(\mathcal{G}),\phi(\mathcal{G}^{\prime})\rangle$, based on these substructures. Weisfeiler-Lehman (WL) graph kernels stand out as one of the most widely used approaches. These methods employ the Weisfeiler-Lehman coloring scheme to iteratively encode graphs, calculating kernel values based on these colorings. The final kernel values are derived through the aggregation of intermediate results. Next, we introduce the WL-subtree kernel and the WLOA kernel in more detail. Let $f^{i}$ be the coloring function at the $i$-th iteration, mapping the colored graph from the previous iteration to a new colored graph. Define $\psi^{i}$ as the function that captures the cumulative coloring effect up to the $i$-th iteration: $\psi^{i}=f^{i}\circ\cdots\circ f^{1}$. Specifically, the WL-subtree kernel computes the kernel value by directly using the label histogram as a feature to compute the dot product, which is expressed as:
$\mathbb{K}^{(h)}_{wl\_subtree}\left(\mathcal{G},\mathcal{G}^{\prime}\right)=\sum_{i=1}^{h}\left\langle\phi(\psi^{i}(\mathcal{G})),\phi(\psi^{i}\left(\mathcal{G}^{\prime})\right)\right\rangle$
The WLOA kernel applies a histogram intersection kernel to match substructures between different graphs, which is formulated as:
$\mathbb{K}^{(h)}_{WLOA}\left(\mathcal{G},\mathcal{G}^{\prime}\right)=\sum_{i=1}^{h}\texttt{histmin}\left\{\phi(\psi^{i}(\mathcal{G})),\phi(\psi^{i}(\mathcal{G}^{\prime}))\right\}\cdot\omega(i)$
where $\omega(i)$ is a nonnegative, monotonically non-decreasing weight function, and $\omega(i)=1$ is commonly used in practice (Kriege et al., 2016; Siglidis et al., 2020). The operator histmin denotes the histogram intersection kernel. It is computed by comparing and summing the smallest matching elements between two sets. For example, consider two sets $S_{1}:\{\mathfrak{a},\mathfrak{a},\mathfrak{b},\mathfrak{b},\mathfrak{c}\}$ and $S_{2}:\{\mathfrak{a},\mathfrak{b},\mathfrak{b},\mathfrak{c},\mathfrak{c}\}$. The $\texttt{histmin}(S_{1},S_{2})$ is calculated by taking the minimum frequency of each distinct element across the two sets, yielding $\min(2,1)+\min(2,2)+\min(1,2)=4$. To ensure that the kernel values fall in the range of 0 to 1, normalization is often applied. The normalized kernel value $\tilde{\mathbb{K}}^{(h)}$ is then expressed as follows:
$\tilde{\mathbb{K}}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})=\frac{\mathbb{K}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})}{\sqrt{\mathbb{K}^{(h)}(\mathcal{G},\mathcal{G})}{\sqrt{\mathbb{K}^{(h)}(\mathcal{G}^{\prime},\mathcal{G}^{\prime})}}}$
3 Consistency Principles
To encode relational structures in GNN learning, we first examine how similarities are represented in graph kernels. In this section, we start by defining a class of graph kernels, i.e., the iterative graph kernels, which encompasses many widely used kernels. Then, we delve into a key property known as the consistency property, which may play an important role in enhancing classification performance. We support this assertion through theoretical analysis, elucidating how different kernels adhere to or deviate from this property, thereby explaining their performance differences.
3.1 Iterative Graph Kernels
In this paper, we are interested in a set of kernels defined as follows:
Definition 3.1
Given a colored graph set $\chi$, a feature map function $\phi:\chi\rightarrow\mathcal{H}_{k}$ (where $\mathcal{H}_{k}$ is a Hilbert space), and a set of coloring functions $\mathcal{F}_{c}=\{f^{0},f^{1},\ldots,f^{i}\}$ on $\chi$ (with $f^{i}:\chi\rightarrow\chi$), we define the set of iterative graph kernels (IGK) as:
$\mathbb{K}_{\mathcal{F}_{c},\phi}(x,y,i)=\langle\phi(f^{i}\circ\cdots f^{1}(x)),\phi(f^{i}\circ\cdots f^{1}(y))\rangle=\langle\psi^{i}(x),\psi^{i}(y)\rangle$
where $x,y\in\chi$ and $\psi^{i}(\cdot)$ represents a composite function given by: $\psi^{i}=f^{i}\circ$ $\cdots\circ f^{1}$, Then the normalized kernel is given by:
$\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)=\frac{\mathbb{K}_{\mathcal{F}_{c},\phi}(x,y,i)}{\sqrt{\mathbb{K}_{\mathcal{F}_{c},\phi}(x,x,i)}\sqrt{\mathbb{K}_{\mathcal{F}_{c},\phi}(y,y,i)}}$
Based on this definition, we can see that graph kernels utilizing the Weisfeiler-Lehman framework, including the WL-subtree kernel (Shervashidze et al., 2011), WLOA (Kriege et al., 2016), and the Wasserstein Weisfeiler-Lehman (WWL) kernel (Togninalli et al., 2019), should be classified as iterative graph kernels. Conversely, the subgraph-pattern approaches, such as the graphlet kernel (Shervashidze et al., 2009) and the shortest-path kernel (Borgwardt and Kriegel, 2005), do not fall into this category.
3.2 Key Properties: Monotonic Decrease & Order Consistency
To effectively capture relational structures, we design the IGKs to progressively differentiate between graphs. With each iteration, additional structural features are considered, enabling the distinction of graphs that may have been indistinguishable in earlier iterations. This implies two properties: (1) the kernel values monotonically decrease with larger iterations, as the similarity between two graphs decreases with the consideration of more features; and (2) the similarity rankings across different iterations should remain consistent, meaning that graphs deemed dissimilar in early iterations should not be considered similar in later iterations. We then formally define these two properties and demonstrate how they can lead to a non-decreasing margin (better performance) in the binary classification task.
Definition 3.2 (Monotonic Decrease)
The normalized iterative graph kernels $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ are said to be monotonically decreasing if and only if:
$\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)\geq\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i+1)\quad\forall x,y\in\chi$
Definition 3.3 (Order Consistency)
The normalized iterative graph kernels $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ are said to preserve order consistency if the similarity ranking remains consistent across different iterations for any pair of graphs, which is defined as:
$\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)>\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,z,i)\Rightarrow\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i+1)\geq\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,z,i+1)\quad\forall x,y,z\in\chi$
Next we show that these two properties can lead to a non-decreasing margin in the binary classification task, which suggests better performance.
Consider a binary graph classification task and assume that the graph representations obtained at any iteration have a uniform norm. This can be achieved by simply normalizing the graph representations at the end of each iteration. That is, for any graph $x$: $\lVert\phi(\psi^{i}(x))\rVert=1$. Then, for any IGK that is monotonically decreasing and preserves the order consistency, the following theorem holds:
Theorem 3.4
Let $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ be a normalized iterative graph kernel that is monotonically decreasing and preserves order consistency. In the binary graph classification task with uniform graph representations, the margin between two classes in the representation space is non-decreasing w.r.t the iteration $i$, where the margin at iteration $i$ is defined as the shortest distance between two graph representations from different classes: margin=$\lVert\psi^{i}(x)-\psi^{i}(y)\rVert$, $\mathcal{Y}(x)\neq\mathcal{Y}(y)$.
Proof sketch. Suppose that at iteration $i$, the margin is defined by a pair of graphs, $x_{1}$ from class 1 and $y_{1}$ from class 2. In the next iteration, $i+1$, the margin is determined by another pair, $x_{2}$ and $y_{2}$. Two possibilities arise: (1) $x_{1}=x_{2}$ and $y_{1}=y_{2}$, or (2) $x_{1}\neq x_{2}$ or $y_{1}\neq y_{2}$.
  1. (1)
    If the pair defining the margin remains unchanged, the margin at iteration $i+1$ can only increase or stay the same, as the kernel is a monotonically decreasing function, indicating a reduction in similarity.
  2. (2)
    We prove case 2 by contradiction, assuming the margin decreases at iteration $i+1$. This would imply that the kernel value for $x_{2}$ and $y_{2}$ increases at iteration $i+1$, which contradicts the fact that the kernel function is monotonically decreasing.
Thus, in both cases, the margin does not decrease as the iterations progress.
We provide the detailed proof in Appendix A.1
3.3 Theoretical Verification with WL-based Kernels
As discussed in Section 3.1, WL-based kernels can be categorized as iterative graph kernels, as they are generated by the coloring functions in an iterative refinement process. Consequently, a natural question arises regarding how various kernels adhere to these properties and whether their adherence reflects their actual performance. We thus investigate two popular WL-based Kernels: the WL-subtree kernel (Shervashidze et al., 2011) and the WLOA kernel (Kriege et al., 2016).
Theorem 3.5
The normalized WL-subtree kernel is neither monotonically decreasing nor does it preserve order consistency.
Theorem 3.6
The normalized WLOA kernel is monotonically decreasing and asymptotically preserves order consistency when $\omega(i)=1$.
Proof sketch. For Theorem 3.5, we illustrate a counterexample, while for Theorem 3.6, we consider two graph pairs where the similarity condition holds:
$\tilde{\mathbb{K}}_{WLOA}^{(h)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)\geq\tilde{\mathbb{K}}_{WLOA}^{(h)}\left(\mathcal{G},\mathcal{G}^{\prime\prime}\right)$
The similarity at the next iteration $h+1$ is scaled by a factor dependent $\omega(i),i=1,\cdots,h+1$. The unnormalized kernel increases monotonically, though with diminishing increments over iterations. Given this, when $\omega(i)=1$ and $h\rightarrow\infty$, we obtain:
$\tilde{\mathbb{K}}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)\geq\tilde{\mathbb{K}}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime\prime}\right)$
We include the complete proof in Appendix A.2 and A.3.
These findings imply that the WLOA kernel better preserves relational structures compared to the WL-subtree kernel, leading to improved classification performance, as supported by the literature (Kriege et al., 2016).
4 Proposed Strategy
Given the analogy between WL-based kernels and GNNs (Shervashidze et al., 2011; Gilmer et al., 2017), and the observation that GNNs often fail to preserve relational structures, we hypothesize that the consistency principle is also beneficial to GNN learning. Thus, we aim to explore how to effectively preserve this principle within the GNN architectures.
Figure 2: Computation of loss. At each layer, pairwise distance matrix $\mathbf{D}$ is calculated using the normalized representations of graphs in a batch. After randomly selecting a reference graph $x_{k}$, the reference probability matrix is computed using the distance matrix from previous layer, where entry $(n,m)$ represents the known probability that the graph $x_{k}$ is more similar to the graph $x_{n}$ than to the graph $x_{m}$. For the distance matrix of current layer, we compute the predicted probability that $x_{k}$ is closer to $x_{n}$ than to $x_{m}$ and form the prediction probability matrix. Consistency loss is computed as the cross-entropy between the predicted and reference probability matrices
4.1 Consistency Loss
Our objective is to enhance graph representation consistency across GNN layers, which has significant potential to preserve the relational structure in the representation space. If we compare $\phi(\psi^{i}(G))$ to the graph representations obtained at the $i$-th layer, preserving the consistency principle is equivalent to preserving the ordering of cosine similarity among the graph representations. However, due to the non-differentiable nature of ranking operations, directly minimizing the ranking differences between consecutive layers is not feasible using gradient-based optimization techniques. Therefore, we aim to optimize pairwise ordering relations instead of the entire ranking list. In this work, our proposed loss employs a probabilistic approach inspired by (Burges et al., 2005). The entire framework is illustrated in Figure 2.
Let $\mathbf{H}^{h}\in\mathbb{R}^{n\times d}$ denote the graph embedding matrix for $n$ examples in a batch, each with a $d$-dimensional feature vector. We first compute the distance matrix $\mathbf{D}^{h}$ for all the graphs in a batch at the $h$-th layer. The entries ${D}_{i,j}^{h}$ of this matrix represent the distance between the representations of the $i$-th and $j$-th graphs, calculated as ${D}_{i,j}^{h}=\texttt{Dist}\left({H}_{x_{i}}^{h},{H}_{x_{j}}^{h}\right)$. Here, we use the cosine distance, the complement of cosine similarity in positive space, expressed as: $1-\frac{{H}_{x_{i}}^{h}\cdot{H}_{x_{j}}^{h}}{\left\|{H}_{x_{i}}^{h}\right\|\cdot\left\|{H}_{x_{j}}^{h}\right\|}$. Considering the distance relationship to an arbitrary graph $x_{k}$ in the batch, the predicted probability ^$\hat{\mathbb{P}}_{n,m|k}^{h}$ that $x_{k}$ is more similar to graph $x_{n}$ than to graph $x_{m}$ at layer $h$ is defined as ^$\hat{\mathbb{P}}_{n,m|k}^{h}\left(\hat{{D}}_{k,n}^{h}<\right.$ $\left.\hat{{D}}_{k,m}^{h}\right)$. This probability score, which ranges from 0 to 1, is formulated using the sigmoid function as follows:
^$\hat{\mathbb{P}}_{n,m|k}^{h}\left(\hat{{D}}_{k,n}^{h}<\hat{{D}}_{k,m}^{h}\right)=\frac{1}{1+\exp\left(\hat{{D}}_{k,m}^{h}-\hat{{D}}_{k,n}^{h}\right)}$
Given the distance matrix from the previous layer ${D}^{h-1}$, the known probability that graph $x_{k}$ is more similar to graph $x_{n}$ than to graph $x_{m}$, denoted as ~$\tilde{\mathbb{P}}_{n,m}^{h-1}(\hat{{D}}_{k,n}^{h-1},\hat{{D}}_{k,m}^{h-1})$, can be formulated as follows:
~$\tilde{\mathbb{P}}_{n,m|k}^{h-1}=\frac{1}{2}\left(1+\operatorname{sign}(\hat{D}_{k,n}^{h-1}-\hat{D}_{k,m}^{h-1})\right)$
We can then minimize the discrepancy between the predicted and the known probability distributions to enhance representation consistency across the layers. Here,we employ the cross-entropy loss to effectively measure the divergence between these two distributions. Specifically, for a pair $\left(x_{n},x_{m}\right)$ centered on $x_{k}$ at layer $h$, the cross-entropy loss can be expressed as:
$\mathcal{L}_{\text{cross-entropy}}\left((x_{n},x_{m})\mid x_{k},h\right)=-\tilde{\mathbb{P}}_{n,m|k}^{h-1}\log\hat{\mathbb{P}}_{n,m|k}^{h}-\left(1-\tilde{\mathbb{P}}_{n,m|k}^{h-1}\right)\log\left(1-\hat{\mathbb{P}}_{n,m|k}^{h}\right)$
Then, the total loss function, which quantifies the consistency of pair-wise distance relations for graph $x_{k}$ at layer $h$, can be formulated as
$\mathcal{L}_{\text{consistency }}(k)=\sum_{n,m}\mathcal{L}((x_{n},x_{m})\mid x_{k},h)$
The overall objective function of the proposed framework can then be formulated as the weighted sum of the original loss and the consistency loss.
$\mathcal{L}_{\text{total }}=\mathcal{L}_{\text{origin}}+\lambda\sum_{i}\mathcal{L}_{\text{consistency }}(i)$
Here, $\lambda$ is a hyperparameter that controls the strength of the consistency constraint.
5 Experiment
In this section, we examine the effectiveness of the proposed consistency loss for the graph classification task. Specifically, we aim to address the following questions: Q1: Does the consistency loss effectively enhance the performance of various GNN backbones in the graph classification task? Q2: How does the consistency loss influence the rank correlation of graph similarities across GNN layers? Q3: How does the consistency loss influence dataset performance across varying levels of complexity, both in structural intricacy and task difficulty?
5.1 Experiment Setup
Dataset
We conduct extensive experiments using the TU Dataset (Morris et al., 2020), the Open Graph Benchmark (OGB) (Hu et al., 2020) and Reddit Threads(Reddit-T) dataset (Bause and Kriege, 2022). The TU Dataset consists of eight graph classification datasets, categorized into three main types: (1) Chem/Bioinformatics datasets, including D&D, NCI1, NCI109, and PROTEINS; (2) Social Network datasets, including IMDB-BINARY,IMDB-MULTI and COLLAB, where a constant feature value of 1 was assigned to all vertices due to the absence of vertex features; and (3) a Computer Vision dataset, COIL-RAG. The OGB datasets include ogbg-molhiv, a molecular property prediction dataset used to determine whether a molecule inhibits HIV replication. Reddit-T dataset is used for classifying threads from Reddit as either discussions or non-discussions, where users are represented as nodes and replies between them as links.
For the TU dataset and Reddit-T, consistent with prior work (Xinyi and Chen, 2019; Xu et al., 2019), we utilize an 8:1:1 ratio for training, validation, and testing sets. For the OGB datasets, we use the official splits provided. Training stops at the highest validation performance, and test accuracy is taken from the corresponding epoch in each fold. Final results are reported as the mean accuracy (except ogbg-molhiv) and standard deviation over $10$ folds. For ogbg-molhiv, we follow the official evaluator and use ROC-AUC as the evaluation metric. Detailed information about these datasets can be found in Appendix B.
Model
We use three widely adopted GNN models as baselines: namely, GCN (Kipf and Welling, 2017), GIN (Xu et al., 2019), and GraphSAGE (Hamilton et al., 2017). We also include two recent GNN models, namely GTransformer (Shi et al., 2021) and GMT (Baek et al., 2021). To ensure a fair comparison, we maintain the same number of layers and layer sizes for both the base models and the models with our proposed consistency loss, ensuring the sharing of the same network architecture. Detailed information about hyperparameter tuning is provided in Appendix C.
5.2 Effectiveness of Consistency Loss
To answer Q1, we present the results for the TU, OGB and Reddit-T datasets in Table 1. As shown in this table, GNN models with the consistency loss yield significant performance improvements over their base models on different datasets. These findings suggest that the consistency framework is a versatile and robust approach for enhancing the predictive capabilities of GNNs in real-world datasets, irrespective of the base model and dataset domain. Notably, the GIN method demonstrates the most significant improvements, achieving enhancements of up to 4.51% on the D&D dataset, 4.32% on the COLLAB dataset, and 3.70% on the IMDB-B dataset. This improvement can be linked to our empirical observation regarding the weak ability of GIN to preserve consistency across layers. In addition, our method demonstrates satisfactory improvements on datasets with numerous classes (e.g., COIL-RAG) and large-scale datasets (e.g., ogbg-molhiv and Reddit-T), indicating that our approach is both flexible and scalable for handling complex and extensive datasets.
Table 1: Classification performance on the TU and OGB datasets, with and without the consistency loss. Highlighted cells indicate instances where the base GNN with the consistency loss outperforms the base GNN alone. The reported values are average accuracy for TU datasets and ROC-AUC for the ogbg-molhiv dataset, including their standard deviations.
NCI1 NCI109 PROTEINS D&D IMDB-B IMDB-M COLLAB COIL-RAG OGB-HIV REDDIT-T #Graphs 4110 4127 1113 1178 1000 1500 5000 3900 41127 203088 Avg. #nodes 29.87 29.68 39.06 284.32 19.77 13.00 74.49 3.01 25.50 23.93 GCN 73.96$\pm$ 2.37 74.04$\pm$ 3.09 73.24$\pm$ 6.93 74.92$\pm$ 2.66 75.40$\pm$ 2.97 55.07$\pm$1.24 81.72$\pm$0.84 91.72$\pm$1.65 72.86$\pm$1.90 76.00$\pm$0.44 +$\mathcal{L}_{\text{consistency}}$ 75.12$\pm$ 1.19 73.25$\pm$ 1.25 75.07$\pm$ 5.05 78.56$\pm$ 3.32 75.85$\pm$ 1.82 56.27$\pm$1.00 83.44$\pm$0.45 93.38$\pm$1.64 73.75$\pm$0.89 77.12$\pm$0.12 GIN 78.13$\pm$ 2.11 76.75$\pm$ 2.91 72.97$\pm$ 4.59 71.10$\pm$ 4.63 70.80$\pm$ 4.07 52.13$\pm$1.42 79.84$\pm$1.05 93.33$\pm$1.48 71.60 $\pm$2.36 77.50$\pm$0.16 +$\mathcal{L}_{\text{consistency}}$ 79.45$\pm$ 1.09 77.46$\pm$ 1.96 74.98$\pm$ 4.57 75.51$\pm$ 2.63 74.50$\pm$ 3.06 53.46$\pm$2.44 84.16$\pm$0.81 94.03$\pm$1.33 74.57$\pm$1.61 77.64$\pm$0.05 GraphSAGE 74.40$\pm$ 1.83 73.17$\pm$ 0.47 74.96$\pm$ 3.14 76.44$\pm$ 4.16 73.90$\pm$ 2.17 51.33$\pm$2.95 78.92$\pm$1.20 89.56$\pm$2.37 77.03$\pm$1.65 76.67$\pm$0.11 +$\mathcal{L}_{\text{consistency}}$ 78.26$\pm$ 1.08 74.10$\pm$ 2.10 76.40$\pm$ 3.12 77.50$\pm$ 3.38 74.75$\pm$ 3.06 54.27$\pm$1.24 82.12$\pm$0.78 92.31$\pm$1.32 78.60$\pm$1.44 77.57$\pm$0.05 GTransformer 75.72$\pm$2.69 74.79$\pm$1.82 73.33$\pm$4.80 75.42$\pm$3.22 72.20$\pm$3.49 53.33$\pm$1.12 80.36$\pm$0.56 83.74$\pm$3.17 76.81$\pm$1.34 76.75$\pm$0.12 +$\mathcal{L}_{\text{consistency}}$ 76.83$\pm$1.36 75.82$\pm$1.53 77.03$\pm$3.79 76.57$\pm$2.54 73.75$\pm$2.56 56.53$\pm$1.54 80.48$\pm$0.47 91.67$\pm$1.88 76.90$\pm$3.25 77.14$\pm$0.06 GMT 75.04$\pm$1.43 73.90$\pm$2.29 72.70$\pm$4.21 72.80$\pm$2.19 79.80$\pm$1.08 54.13$\pm$2.90 80.36$\pm$1.15 90.85$\pm$1.91 74.86$\pm$2.26 72.06$\pm$10.15 +$\mathcal{L}_{\text{consistency}}$ 75.52$\pm$1.07 75.20$\pm$0.95 74.86$\pm$2.03 73.14$\pm$2.28 79.60$\pm$1.91 54.80$\pm$1.42 82.80$\pm$0.61 92.00$\pm$1.43 76.00$\pm$1.99 77.19$\pm$0.14
Furthermore, we analyze the complexity and scalability of our method on the TU Dataset, with additional details provided in Appendices D.1 and D.2. Additionally, we demonstrate the method’s potential to enhance performance significantly, even with a marginal increase in computational cost, as illustrated in Appendix E.
5.3 Effect of the Consistency Loss on Rank Correlation
Table 2: Spearman correlation was computed for graph representations from consecutive layers on the TU datasets, both with and without consistency loss. Values with higher rank correlation are highlighted in grey. The consistency loss can enhance the rank correlation of graph similarities.
NCI1 NCI109 PROTEINS D&D IMDB-B GCN 0.753 0.920 0.584 0.709 0.846   +$\mathcal{L}_{\text{consistency}}$ 0.859 0.958 0.946 0.896 0.907 GIN 0.666 0.674 0.741 0.721 0.598   +$\mathcal{L}_{\text{consistency}}$ 0.877 0.821 0.904 0.847 0.816 GraphSAGE 0.903 0.504 0.845 0.741 0.806   +$\mathcal{L}_{\text{consistency}}$ 0.911 0.709 0.916 0.872 0.933 GTransformer 0.829 0.817 0.867 0.865 0.884   +$\mathcal{L}_{\text{consistency}}$ 0.863 0.883 0.915 0.880 0.917 GMT 0.872 0.887 0.980 0.826 0.893   +$\mathcal{L}_{\text{consistency}}$ 0.906 0.908 0.983 0.856 0.908
To answer Q2, we compare the consistency of graph representations across layers with and without the proposed consistency loss. Specifically, we use the Spearman’s rank correlation coefficient, a widely accepted method for computing correlation between ranked variables, to quantitatively measure the consistency of graph similarities across layers. For a fair comparison, we construct a distance matrix $D^{h}$ for all test data at each layer $h$, where each row $D_{x_{i};}^{h}$ represents the distances from graph $x_{i}$ to all other graphs. We then compute the rank correlation between $D_{x_{i},:}^{h}$ and $D_{x_{i};}^{h+1}$ for each graph $x_{i}$.
We average the correlation values for all graphs to obtain the overall correlation for layer $h$. Then, we compute the mean of these values across layers, enabling a global comparison of relational consistency throughout the model and dataset. All results were averaged over 5 repeated experiments with same training setting.
We present our results on a series of datasets from the TU Dataset in Table 2. As shown in the table, it is evident that the representation space becomes more consistent with our proposed consistency loss. For example, a significant enhancement is observed for the GIN model. Another notable point is the result for the GCN model on the NCI109 dataset and for the GMT model on the IMDB-B dataset. We find that the correlation is already fairly high even without the implementation of $\mathcal{L}_{\text{consistency}}$, resulting in minimal correlation improvements with our method. This phenomenon provides a plausible explanation for why our method is not effective in these two cases.
5.4 Study on Task Complexity
To address Q3 and further evaluate the performance of our method across different scenarios, we extended our study by conducting experiments on graph datasets with increasing task and structural complexity.
Increasing Task Complexity
We increase the task complexity by expanding the number of classes that the model needs to classify. To assess the effect of increased class complexity on our method’s performance, we sampled subsets from the REDDIT-MULTI 5K dataset (Yanardag and Vishwanathan, 2015) with a progressively greater number of classes, which originally consists of five classes. Specifically, we randomly sampled between 2 and 4 classes to construct new datasets from the original dataset and conducted classification tasks using both GCN and GCN with $\mathcal{L}_{\text{consistency}}$ on these newly constructed datasets. We report the mean test accuracy over five experiments for each subset, as presented in Table.3.
Table 3: Performance comparison across different subsets and the full set.
Subset1 Subset2 Subset3 Fullset
(2 classes) (3 classes) (4 classes) (5 classes)
GCN 79.50 67.13 50.30 53.80
GCN+$\mathcal{L}_{\text{consistency}}$ 81.10 68.00 57.15 57.12
The results demonstrate that the effectiveness of our method remains robust, even as the number of classes increases. In fact, it may provide greater advantages when applied to multi-class classification tasks. This resilience likely stems from our method’s focus on identifying relational structures in the intermediate representations of GNN models, rather than relying heavily on label information. This approach helps mitigate the impact of potential label noise in the original data. These findings align with the noticeable performance improvements observed in both binary and multi-class classification tasks, as shown in Table 1.
Increasing Structural Complexity
We assessed the impact of structural complexity by partitioning the IMDB-B dataset into three subsets with progressively increasing graph densities. Graph density, denoted as $d=\frac{2M}{N(N-1)}$, where $N$ is the number of nodes and $M$ is the number of edges in graph $\mathcal{G}$, was used as the criterion for creating these subsets. The dataset was divided into three groups: (small) for graphs with densities below the 33rd percentile, (median) for densities between the 33rd and 67th percentiles, and (large) for graphs with densities above the 67th percentile. We applied both GCN and GCN+$\mathcal{L}_{\text{consistency}}$ models to these subsets, and the results are summarized in Table 4.
Table 4: Performance comparison on IMDB-B datasets of different densities.
IMDB-B IMDB-B IMDB-B
(small) (medium) (large)
GCN 77.58±$\pm 4.11$ 66.25±$\pm 5.38$ 67.61±$\pm 6.21$
GCN+$\mathcal{L}_{\text{consistency}}$ 84.24±$\pm 4.85$ 69.06±$\pm 4.06$ 71.43±$\pm 4.43$
The above results show that the GCN model, enhanced with the $\mathcal{L}_{\text{consistency}}$ loss function, consistently outperforms the original version across different structural complexity groups, demonstrating the robustness and effectiveness of the proposed method.
Furthermore, we evaluate the method’s efficiency under various complexity scenarios, as detailed in the Appendix D.3.
6 Related Work
Graph Distance and Similarity
Measuring distances or similarities between graphs is a fundamental problem in graph learning. Graph kernels, which define graph similarity, have gained significant attention. Most graph kernels use the $\mathcal{R}$-Convolution framework (Haussler et al., 1999) to compare substructure similarities. A trailblazing kernel by (Kashima et al., 2003) used node and edge attributes to generate label sequences through a random walk. The WL-subtree kernel (Shervashidze et al., 2011) generates graph-level features by summing node representation contributions. Recent works align matching substructures between graphs. For instance, Kriege et al. (2016) proposed a discrete optimal assignment kernel based on vertex kernels from WL labels. Togninalli et al. (2019) extended this to include fractional assignments using the Wasserstein distance. Additionally, measuring graph distances is also a prevalent problem. Vayer et al. (2019) combined the Wasserstein and Gromov-Wasserstein distances (Villani and Society, 2003; Mémoli, 2011). Chen et al. (2022) proposed a polynomial-time WL distance for labeled Markov chains, treating labeled graphs as a special case.
Bridging Graph Kernels and GNNs
Many studies have explored the connection between graph kernels and GNNs, attempting to integrate them into a unified framework. Certain approaches focus on leveraging GNN architecture to design novel kernels. For instance, Mairal et al. (2014) presents neural network architectures that learn graph representations within the Reproducing Kernel Hilbert Space (RKHS) of graph kernels. Similarly, Du et al. (2019) proposed a graph kernel equivalent to infinitely wide GNNs, which can be trained using gradient descent. Conversely, other studies have incorporated kernel methods directly into GNNs. For example, Nikolentzos and Vazirgiannis (2020) utilize graph kernels as convolutional filters within GNN architectures. Additionally, Lee et al. (2024) proposes a novel Kernel Convolution Network that employs the random walk kernel as the core mechanism for learning descriptive graph features. Instead of applying specific kernel patterns as mentioned in previous work, we introduce a general method for GNNs to capture consistent similarity relationships, thereby enhancing classification performance.
7 Conclusion
In this paper, we study a class of graph kernels and introduce the concept of consistency property in graph classification tasks. We theoretically prove that this property leads to a more structure-aware representation space for classification using kernel methods. Based on this analysis, we extend this principle to enhance GNN models. We propose a novel, model-agnostic consistency learning framework for GNNs that enables them to capture relational structures in the graph representation space. Experiments show that our proposed method universally enhances the performance of backbone networks on graph classification benchmarks, providing new insights into bridging the gap between traditional kernel methods and GNN models.
Appendix A Theoretical Verification with WL-based Kernels
A.1 Non-decreasing Margin
Suppose that at iteration $i$, graph $x_{1}$ in class 1 and graph $y_{1}$ in class 2 decide the margin and at iteration $i+1$, graph $x_{2}$ in class 1 and graph $y_{2}$ in class 2 decide the margin.
There are two scenarios: (1) $x_{1}=x_{2}$ and $y_{1}=y_{2}$, (2) $x_{1}\neq x_{2}$ or $y_{1}\neq y_{2}$. For case 1, margin at iteration $i$ is given by:
$\begin{split}\lVert\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1}))\rVert&=\sqrt{(\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1})))^{T}(\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1})))}\\ &=\sqrt{2-2\langle\phi(\psi^{i}(x_{1})),\phi(\psi^{i}(y_{1}))\rangle}=\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{1},y_{1},i)}.\end{split}$ (1)
Since $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ is monotonically decreasing, we have:
$\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{1},y_{1},i)}\leq\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{1},y_{1},i+1)}=\lVert\phi(\psi^{i+1}(x_{1}))-\phi(\psi^{i+1}(y_{1}))\rVert.$ (2)
This suggests that the margin does not decrease.
For case 2, we prove by contradiction and first assume that the margin decreases. Since the margin decreases, we have:
$\lVert\phi(\psi^{i+1}(x_{2}))-\phi(\psi^{i+1}(y_{2}))\rVert<\lVert\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1}))\rVert.$ (3)
At iteration $i$, since $x_{1}$ and $y_{1}$ decide the margin, we have:
$\lVert\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1}))\rVert\leq\lVert\phi(\psi^{i}(x_{2}))-\phi(\psi^{i}(y_{2}))\rVert.$ (4)
Combining Equation 3 and Equation 4, we have:
$\lVert\phi(\psi^{i+1}(x_{2}))-\phi(\psi^{i+1}(y_{2}))\rVert<\lVert\phi(\psi^{i}(x_{2}))-\phi(\psi^{i}(y_{2}))\rVert.$ (5)
This is equivalent to:
$\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{2},y_{2},i+1)}<\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{2},y_{2},i)},$ (6)
[…]
A.2 Proof of Theorem 3.5
Following Equation 2.4, the normalized WL-subtree kernel after $h$ iterations can be expressed as:
$\tilde{\mathbb{K}}^{(h)}_{wl\_subtree}\left(\mathcal{G},\mathcal{G}^{\prime}\right)=\frac{\sum_{i=1}^{h}\left\langle\phi(\psi^{i}(\mathcal{G})),\phi(\psi^{i}\left(\mathcal{G}^{\prime})\right)\right\rangle}{\sqrt{\sum_{i=1}^{h}\left\langle\phi(\psi^{i}(\mathcal{G})),\phi(\psi^{i}\left(\mathcal{G})\right)\right\rangle}\sqrt{\sum_{i=1}^{h}\left\langle\phi(\psi^{i}(\mathcal{G}^{\prime})),\phi(\psi^{i}\left(\mathcal{G}^{\prime})\right)\right\rangle}}$
Next, we show by counterexample that the WL-subtree kernel may not preserve relational structure.
For example, consider two graphs, $\mathcal{G}$ and $\mathcal{G}^{\prime}$, in the $h$-th iteration. The color (label) set of the nodes in these graphs is denoted as $\left\{C_{a},C_{b},C_{c},\ldots\right\}$. Graph $\mathcal{G}$ contains 200 nodes labeled $C_{a}$ and 4 nodes labeled $C_{b}$, whereas graph $\mathcal{G}^{\prime}$ contains 4 nodes labeled $C_{a}$ and 200 nodes labeled $C_{b}$. The similarity is computed using the vectors $\phi\left(\psi^{i}(\mathcal{G})\right)=[200,4]$ and $\phi\left(\psi^{i}\left(\mathcal{G}^{\prime}\right)\right)=[4,200]$. Consequently, the resulting similarity $\tilde{\mathbb{K}}^{(h)}_{wl\_subtree}$ is 0.0400.
In the next iteration, for graph $\mathcal{G}$, assume that half of the nodes currently labeled $C_{a}$ are relabeled to $C_{c}$, while the other half are relabeled to $C_{d}$. Additionally, all nodes labeled $C_{b}$ are relabeled to $C_{e}$. As a result, the updated histogram vector for $\mathcal{G}$ becomes $\phi\left(\psi^{i+1}(\mathcal{G})\right)=[100,100,4]$. Similarly, in graph $\mathcal{G}^{\prime}$, half of the nodes labeled $C_{a}$ are relabeled to $C_{c}$ and the remaining half to $C_{d}$, with all nodes labeled $C_{b}$ relabeled to $C_{e}$, resulting in a histogram vector of $\phi\left(\psi^{i+1}(\mathcal{G}^{\prime})\right)=[2,2,200]$. Here, the normalized similarity $\tilde{\mathbb{K}}^{(h)}_{wl\_subtree}$ is 0.0404, which exceeds $\tilde{\mathbb{K}}^{(h)}_{wl\_subtree}$.
[…]
A.3 Proof of Theorem 3.6
A.3.1 Monotonic Decrease in WLOA Kernel
The WLOA kernel after $h$ iterations can be computed as follows:
$\mathbb{K}^{(h)}_{WLOA}\left(\mathcal{G},\mathcal{G}^{\prime}\right)=\sum_{i=1}^{h}\texttt{histmin}\left\{\phi(\psi^{i}(\mathcal{G})),\phi(\psi^{i}(\mathcal{G}^{\prime}))\right\}\cdot\omega(i)$
where $\omega(i)$ is a monotonically increasing function. The normalized kernel is given by:
$\tilde{\mathbb{K}}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})=\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})}{\sqrt{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G})}{\sqrt{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G}^{\prime},\mathcal{G}^{\prime})}}}$
At any iteration $i$, the expression $\texttt{histmin}\left\{\phi(\psi^{i}(\mathcal{G})),\phi(\psi^{i}(\mathcal{G}))\right\}\cdot\omega(i)$ equals to $\omega(i)|\mathcal{V}|$. Consequently, for the $h$-th iteration, this implies:
$\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G})=\sum_{i=1}^{h}\omega(i)|\mathcal{V}|$
Thus, the normalized kernel value for the graph pair $\left(\mathcal{G},\mathcal{G}^{\prime}\right)$ can be expressed as:
$\tilde{\mathbb{K}}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})=\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})}{\sqrt{|\mathcal{V}||\mathcal{V}^{\prime}|}\sum_{h}^{i=1}\omega(i)}$
[…]
$\tilde{\mathbb{K}}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})-\tilde{\mathbb{K}}_{WLOA}^{(h+1)}(\mathcal{G},\mathcal{G}^{\prime})$
$=$ $\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})}{\sqrt{|\mathcal{V}||\mathcal{V}^{\prime}|}\sum_{i=1}^{h}\omega(i)}-\frac{\mathbb{K}_{WLOA}^{(h+1)}(\mathcal{G},\mathcal{G}^{\prime})}{\sqrt{|\mathcal{V}||\mathcal{V}^{\prime}|}\sum_{i=1}^{h+1}\omega(i)}$
$=$ $\frac{1}{\sqrt{|\mathcal{V}||\mathcal{V}^{\prime}|}}\cdot(\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})}{\sum_{i=1}^{h}\omega(i)}-\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})+\texttt{histmin}\{\phi(\psi^{h+1}(\mathcal{G})),\phi(\psi^{h+1}(\mathcal{G}^{\prime}))\omega(h+1)\}}{\sum_{i=1}^{h+1}\omega(i)})$
::=:= $D$
Considering the hierarchical structure of the refinement processes, one can deduce the following inequality:
$\texttt{histmin}\{\phi(\psi^{i}(\mathcal{G})),\phi(\psi^{i}(\mathcal{G}^{\prime}))\}\leq\texttt{histmin}\{\phi(\psi^{i-1}(\mathcal{G})),\phi(\psi^{i-1}(\mathcal{G}^{\prime}))\}\leq\cdots\texttt{histmin}\{\phi(\psi^{1}(\mathcal{G})),\phi(\psi^{1}(\mathcal{G}^{\prime}))\}$
Thus, for the kernel at the $(h+1)$-th iteration, it follows that:
$\texttt{histmin}\{\phi(\psi^{h+1}(\mathcal{G})),\phi(\psi^{h+1}(\mathcal{G}^{\prime}))\}\leq\frac{\sum_{i=1}^{h}\texttt{histmin}\{\phi(\psi^{i}(\mathcal{G})),\phi(\psi^{i}(\mathcal{G}^{\prime}))\cdot\omega(i)\}}{\sum_{i=1}^{h}\omega(i)}=\frac{\mathbb{K}_{WLOA}^{(h)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)}{\sum_{i=1}^{h}\omega(i)}$
Given this inequality, we get
$D$ $\geq\frac{1}{\sqrt{|\mathcal{V}||\mathcal{V}^{\prime}|}}\cdot(\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})}{\sum_{i=1}^{h}\omega(i)}-\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})+\frac{\omega(h+1)}{\sum_{i=1}^{h}\omega(i)}\mathbb{K}_{WLOA}^{(h)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)}{\sum_{i=1}^{h+1}\omega(i)})$ (7)
$=\frac{1}{\sqrt{|\mathcal{V}||\mathcal{V}^{\prime}|}}\cdot(\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})}{\sum_{i=1}^{h}\omega(i)}-\frac{\sum_{i=1}^{h}\omega(i)\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})}{\sum_{i=1}^{h}\omega(i)}+\frac{\omega(h+1)}{\sum_{i=1}^{h}\omega(i)}\mathbb{K}_{WLOA}^{(h)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)}{\sum_{i=1}^{h+1}\omega(i)})$
$=\frac{1}{\sqrt{|\mathcal{V}||\mathcal{V}^{\prime}|}}\cdot\frac{\mathbb{K}_{WLOA}^{(h)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)}{\sum_{i=1}^{h}\omega(i)}\cdot(1-\frac{\sum_{i=1}^{h}\omega(i)+\omega(h+1)}{\sum_{i=1}^{h}\omega(i)+\omega(h+1)})$
$=0$
Therefore, we get $D\geq 0$. The value of Equation 7 is non-negative, indicating that WLOA is monotonically decreasing.
A.3.2 Order Consistency in WLOA Kernel
Assume two graph pairs satisfy the following relation:
$\tilde{\mathbb{K}}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})\geq\tilde{\mathbb{K}}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime\prime})$
The normalized WLOA kernel for a graph pair $\left(\mathcal{G},\mathcal{G}^{\prime}\right)$ at iteration $h+1$ can be expressed as:
$\tilde{\mathbb{K}}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)=\frac{\mathbb{K}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)}{\sqrt{|\mathcal{V}|\left|\mathcal{V}^{\prime}\right|}\sum_{i=1}^{h+1}\omega(i)}=\frac{\sum_{i=1}^{h}\omega(i)}{\sum_{i=1}^{h+1}\omega(i)}\cdot\frac{\mathbb{K}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)}{\sqrt{|\mathcal{V}|\left|\mathcal{V}^{\prime}\right|}\sum_{i=1}^{h}\omega(i)}$
$\mathbb{K}_{WLOA}^{(h)}$ is monotonically increasing with the iteration number $h$, as it follows from: $\mathbb{K}_{WLOA}^{(h+1)}=\mathbb{K}_{WLOA}^{(h)}+\texttt{histmin}\left\{\phi(\psi^{h+1}(\mathcal{G})),\phi(\psi^{h+1}(\mathcal{G}^{\prime}))\right\}\cdot\omega(h+1)$. Thus, we have:
$\tilde{\mathbb{K}}_{WLOA}^{(h+1)}(\mathcal{G},\mathcal{G}^{\prime})$ $\geq\frac{\sum_{i=1}^{h}\omega(i)}{\sum_{i=1}^{h+1}\omega(i)}\cdot\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime})}{\sqrt{|\mathcal{V}||V^{\prime}|}\sum_{i=1}^{h}\omega(i)}$
$\geq\frac{\sum_{i=1}^{h}\omega(i)}{\sum_{i=1}^{h+1}\omega(i)}\cdot\frac{\mathbb{K}_{WLOA}^{(h)}(\mathcal{G},\mathcal{G}^{\prime\prime})}{\sqrt{|\mathcal{V}||\mathcal{V}^{\prime\prime}|}\sum_{i=1}^{h}\omega(i)}$
$=\frac{\sum_{i=1}^{h}\omega(i)}{\sum_{i=1}^{h+1}\omega(i)}\cdot\tilde{\mathbb{K}}_{WLOA}^{(h)}\left(\mathcal{G},\mathcal{G}^{\prime\prime}\right)$
Given the monotonic decrease property of $\tilde{\mathbb{K}}_{WLOA}^{(h)}\left(\mathcal{G},\mathcal{G}^{\prime\prime}\right)$,we have:
$\tilde{\mathbb{K}}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime\prime}\right)\leq\tilde{\mathbb{K}}_{WLOA}^{(h)}\left(\mathcal{G},\mathcal{G}^{\prime\prime}\right)$
Thus, we can conclude:
$\tilde{\mathbb{K}}_{WLOA}^{(h+1)}(\mathcal{G},\mathcal{G}^{\prime})$ $\geq\frac{\sum_{i=1}^{h}\omega(i)}{\sum_{i=1}^{h+1}\omega(i)}\tilde{\mathbb{K}}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime\prime}\right)$
$=\left(1-\frac{\omega(h+1)}{\sum_{i=1}^{h+1}\omega(i)}\right)\tilde{\mathbb{K}}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime\prime}\right)$
[…]
Appendix B Dataset
In this section, we provide the statistics of the datasets used, as shown in Table 5.
Table 5: Dataset statistics.
Dataset Task Description # Class # Size Ave.Nodes Ave.Edges Node Label.
ogbg-molhiv Molecular property prediction 2 41127 25.5 54.1 +
PROTEINS Enzyme classification 2 1113 39.06 72.82 +
D&D Enzyme classification 2 1178 284.32 715.66 +
NCI1 Molecular classification 2 4110 29.87 32.30 +
NCI109 Molecular classification 2 4127 29.68 32.13 +
IMDB-B Movie venue categorization 2 1000 19.77 96.53 -
IMDB-M Movie venue categorization 3 1500 13.00 65.94 -
COLLAB Collaboration classification 3 5000 74.49 2457.78 -
COIL-RAG Computer Vision 100 3900 3.01 3.02 -
Reddit-T Reddit Thread Classification 2 203088 23.93 24.99 -
Appendix C Detailed Set-Up
TU Dataset
We restrict the hyperparameters and ensure the same architecture is used for both the base and enhanced models on the same dataset for a fair comparison. Specifically, we fixed the hidden size to 32, the number of layers to 3, and used a global mean pooling layer to generate the graph representations. The Adam optimizer was used for optimization. During the training process, we tuned both the base and enhanced models with the same search ranges: batch size {64, 128}, dropout rate {0, 0.3} and learning rate {0.0001, 0.001, 0.01}.,The only additional parameter for the enhanced model is the regularization term, which ranges from {0.1, 0.5, 1, 10}.
OGB
Considering the complexity of the OGB dataset, we slightly expand the hyperparameter search range. Initially, we train the model with consistency loss, exploring hidden sizes {32, 64, 128, 256}, batch sizes {64, 128, 256}, and the number of layers {3, 4}, while keeping other parameters consistent with our experiments on the TU dataset. Subsequently, we fix the hidden size and the number of layers to ensure an identical network structure. We then repeat the parameter search process and employ the optimal settings for testing on the base model. All experiments on the OGB dataset are repeated 5 times to calculate the mean and standard deviation.
Reddit-T
Given that the Reddit-T and OGB datasets are of comparable size, we employ similar experimental settings for training models on these datasets. The architecture is kept fixed, and we search for optimal hyperparameters for both the base model and the models incorporating consistency loss. Each model is trained and evaluated over 5 independent runs, with the mean and standard deviation of the results recorded for performance comparison.
Appendix D Complexity& Scalability
D.1 Time Complexity
We present the time complexity analysis for our proposed consistency loss. The loss computation involves the compuation of pairwise similarities of graphs in a batch, resulting in a computation complexity of $O\left(\text{batchsize}\cdot\frac{\text{batchsize}-1}{2}\right)=O(\text{batchsize}^{2})$. Given that there are $\frac{\text{datasetsize}}{\text{batchsize}}$ batches in each training epoch and that the similarities are computed between consecutive layers, the total complexity is:
$O(\text{ loss })$ $=O\left\{\text{ batchsize}^{2}\times(\text{ layernum }-1)\times\frac{\text{ datasetsize }}{\text{ batchsize }}\right\}$
$=O(\text{ datasetsize}\times\text{batchsize}\times\text{layernum }).$
This analysis shows that the time required to compute consistency loss scales linearly with dataset size, batch size, and the number of layers. It is important to note that the training time for baseline models also scales linearly with dataset size.
Figure 3: Training Cost Escalates Linearly with Dataset Size Increase
Since batch size and the number of layers are generally small compared to dataset size, our experiments primarily focus on how dataset size affects training time. We evaluate the training time of several models—GCN, GIN, GraphSAGE, GTransformer, and GMT—each enhanced with our consistency loss. This evaluation is conducted on different subsets of the ogbg-molhiv dataset, with subset sizes adjusted by varying the sampling rates. The training time, measured in seconds, are presented in Figure 3 . As shown, our findings confirm that training time increases linearly with dataset size, indicating that our method maintains training efficiency comparable to baselines without adding significant time burdens.
Furthermore, we empirically measure the training time for both the baseline models and our proposed methods. Each model comprises three layers and is trained on the ogbg-molhiv dataset (40,000+ graphs) for 100 epochs. We calculate the average training time per epoch in seconds and present the results in Table 6, showing that while the inclusion of the consistency loss slightly increases the training time, the impact is minimal.
Table 6: Average training time per epoch for different models on the ogbg-molhiv dataset, measured in seconds.
GMT GTransformer GIN GCN GraphSAGE
GCN 8.380 4.937 4.318 4.221 3.952
GCN+$\mathcal{L}_{\text{consistency}}$ 8.861 6.358 5.529 5.382 5.252
D.2 Space Complexity
Next, we present the space complexity analysis for our consistency loss. At each iteration, the loss function requires storing two pairwise similarity matrices corresponding to two consecutive layers, which is given by:
$O(\text{ loss })=O(\text{batchsize}^{2})$
Since we use stochastic gradient descent, similarity matrices are not retained for the next iteration. The consistency loss requires significantly less space than node embeddings, making the additional space requirement minimal. Table 7 shows the peak memory usage in megabytes (MB) for different models when training on the ogbg-molhiv dataset, illustrating that the space costs are negligible.
Table 7: Peak memory usage for different models on the ogbg-molhiv dataset, measured in megabytes.
GMT GTransformer GIN GCN GraphSAGE
GCN 1334.0 1267.8 1291.3 1274.2 1288.4
GCN+$\mathcal{L}_{\text{consistency}}$ 1370.0 1330.6 1338.9 1320.1 1321.3
Cost Increase (%) 2.70 4.96 3.68 3.60 2.55
D.3 Efficiency on Different Task and Structural Complexities
Task Complexity
We measured the runtime of the models on different subsets to evaluate how task complexity, in terms of the number of classes, influences the efficiency of the proposed method. The results are presented in Table 8. As demonstrated, the additional computational time remains minimal even with an increasing number of classes, suggesting that the method scales effectively with growing class complexity.
Table 8: Average training time per epoch on REDDIT subsets with varying class complexity, measured in seconds
Subset1 Subset2 Subset3 Fullset
(2 classes) (3 classes) (4 classes) (5 classes)
GCN $0.203$ $0.345$ $0.408$ $0.493$
GCN+$\mathcal{L}_{\text{consistency}}$ $0.227$ $0.355$ $0.430$ $0.557$
Structure Complexity
We also conducted experiments to assess the training costs on datasets with varying structural complexities when introducing the $\mathcal{L}_{\text{consistency}}$ The results, summarized below in Table 9, show that the additional training cost remains minimal across datasets with different structures. This demonstrates the broad applicability of the proposed method, regardless of structural complexity.
Table 9: Average training time per epoch for subsets of varying structural complexity from IMDB-B, measured in seconds.
IMDB-B IMDB-B IMDB-B
(small) (medium) (large)
GCN $0.0308$ $0.0311$ $0.0321$
GCN+$\mathcal{L}_{\text{consistency}}$ $0.0371$ $0.0378$ $0.0392$
Appendix E Efficient Consistent Learning
To further minimize the overhead of our proposed consistency loss, we examined a scenario where the consistency loss, denoted as $\mathcal{L}_{FL}$, was only applied to the first and last layers.
Building upon the experimental setup described in Section 5, we conducted experiments using various backbone models. The results are summarized in Table 10. The penultimate column of this table highlights the performance gains achieved by applying the consistency loss across all layers, while the final column demonstrates the improvements observed when the consistency loss is only applied to the first and last layers.
Table 10: Graph classification performance with improvements of $\mathcal{L}_{ALL}$ and $\mathcal{L}_{FL}$ over base models.
NCI1 NCI109 PROTEINS DD IMDB-B OGB-HIV $\mathcal{L}_{ALL}$uparrow $\mathcal{L}_{FL}$uparrow GCN+$\mathcal{L}_{FL}$ 75.96±$\pm 0.89$ 74.67±$\pm 1.11$ 72.97±$\pm 2.85$ 76.27±$\pm 1.69$ 74.6±$\pm 1.85$ 74.44±$\pm 1.42$ +5.49 +7.08 GIN+$\mathcal{L}_{FL}$ 79.08±$\pm 1.21$ 77.0±$\pm 2.01$ 73.15±$\pm 2.76$ 74.07±$\pm 1.38$ 74.8±$\pm 4.66$ 74.2±$\pm 1.62$ +10.95 +15.12 GraphSAGE+$\mathcal{L}_{FL}$ 78.88±$\pm 2.01$ 74.24±$\pm 1.21$ 75.32±$\pm 2.46$ 73.90±$\pm 2.03$ 76.6±$\pm 1.96$ 80.06±$\pm 1.21$ +9.10 +9.71 GTransformer+$\mathcal{L}_{FL}$ 76.79±$\pm 1.24$ 74.38±$\pm 0.49$ 73.69±$\pm 2.09$ 75.08±$\pm 1.57$ 76.80±$\pm 1.60$ 80.53±$\pm 0.73$ +9.00 +8.63 GMT+$\mathcal{L}_{FL}$ 76.40±$\pm 1.00$ 75.64±$\pm 0.77$ 72.25±$\pm 3.96$ 73.39±$\pm 2.18$ 76.00±$\pm 1.36$ 81.05±$\pm 1.29$ +6.23 +5.24
Notably, applying the consistency loss only to the first and last layers achieves performance comparable to that of applying it across all layers, with both configurations yielding substantial improvements over the original model. This finding suggests that our proposed approach can be accelerated with minimal additional computational cost while still enhancing performance, thereby validating the effectiveness of the consistent learning principle.
Appendix F Similarity/Difference with Contrastive learning
In this section, we discuss the similarities and differences between our method and graph contrastive learning. Graph Contrastive Learning (GCL) is a self-supervised technique for graph data that emphasizes instance discrimination [Lin et al., 2023, Zhu et al., 2021]. A typical GCL framework generates multiple graph views via augmentations and contrasts positive samples (similar instances) with negative samples (dissimilar instances). This approach facilitates effective representation learning by capturing relationships between views, ensuring positive pairs remain close in the embedding space while distinctly separating negative pairs.
While both GCL and our method leverage graph similarity, our approach focuses on maintaining consistency across layers, rather than solely capturing similarities as in contrastive learning. To demonstrate this, we integrated the GraphCL technique [You et al., 2020] into a GCN model (GCN+CL) and assessed its performance and layer consistency across various datasets. The results, detailed in Tables 11 and 12, use classification accuracy and Spearman rank correlation to measure performance and consistency, respectively.
Table 11: Graph classification accuracy of GCN with contrastive learning applied across various datasets.
NCI1 NCI109 PROTEINS D&D IMDB-B
GCN+ CL 74.06 ±$\pm 1.91$ 73.14 ±$\pm 1.90$ 72.50 ±$\pm 2.73$ 75.80 ±$\pm 2.09$ 75.80 ±$\pm 1.90$
GCN+$\mathcal{L}_{\text{consistency}}$ 75.12 ±$\pm 1.19$ 73.25 ±$\pm 1.25$ 75.07 ±$\pm 5.05$ 78.56 ±$\pm 3.32$ 75.85 ±$\pm 1.82$
Table 12: Spearman correlation for graph representations from consecutive layers.
NCI1 NCI109 PROTEINS D&D IMDB-B
GCN+ CL 0.835 0.717 0.851 0.717 0.810
GCN+$\mathcal{L}_{\text{consistency}}$ 0.859 0.958 0.946 0.896 0.907
As demonstrated by the results, our method consistently outperforms GCN+CL in both graph classification performance and in enhancing similarity consistency across layers. This underscores the significant differences between our approach and regular GCL methods.
Appendix G Boarder Impact
This paper aims to advance the field of graph learning by proposing a model-agnostic consistency learning framework. Our framework can be plugged into and improve current methods for graph classification tasks. This has potential benefits in sectors such as chemistry, bioinformatics and social analysis, where graph classification is widely used. Additionally, we do not foresee any direct negative societal or ethical consequences stemming from our work.
Appendix H Limitation
One limitation of our work is that the method involves additional computational costs, especially during large batch training processes. To extend our framework, sampling methodologies on data or layers can be applied during the consistency-preserving training process. By selectively sampling data points or specific layers, we can reduce the computational burden while still maintaining the effectiveness of the cross-layer consistency loss, making the framework more scalable and applicable to larger datasets.

### v2

@@
 Graph kernel methods, on the other hand, are designed to capture similarities between graphs and utilize these similarities for classification tasks.
-For instance, subgraph-pattern approaches (Shervashidze et al., 2009; Costa and Grave, 2010; Kriege et al., 2020) compare graphs by counting the occurrences of fixed-size subgraph motifs.
+For instance, subgraph-pattern approaches (Shervashidze et al., 2009; Kriege et al., 2020) compare graphs by counting the occurrences of fixed-size subgraph motifs.
 Other methods compare sequences of vertices or edges encountered during graph traversals (Borgwardt and Kriegel, 2005; Kashima et al., 2003; Zhang et al., 2018).
@@
 Within this framework, we define the consistency property, which ensures that similar graphs remain similar in subsequent iterations.
-Our analysis demonstrates that kernels with this property yield better classification performance.
-Furthermore, we find that this property sheds light on why the WLOA kernel outperforms the WL-subtree kernel.
+Our analysis demonstrates that kernels with this property ensure better generalization ability, leading to improved classification performance.Furthermore, we find that this property sheds light on why the WLOA kernel outperforms the WL-subtree kernel.
 The WLOA kernel asymptotically demonstrates consistency as the iteration goes to infinity, whereas the WL-subtree kernel does not exhibit this behavior.
@@
 This implies two properties: (1) the kernel values monotonically decrease with larger iterations, as the similarity between two graphs decreases with the consideration of more features; and (2) the similarity rankings across different iterations should remain consistent, meaning that graphs deemed dissimilar in early iterations should not be considered similar in later iterations.
-We then formally define these two properties and demonstrate how they can lead to a non-decreasing margin (better performance) in the binary classification task.
+We then provide formal definitions for these two properties and prove how they can improve generalization in the binary classification task.
 Definition 3.2 (Monotonic Decrease)
@@
 $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)>\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,z,i)\Rightarrow\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i+1)\geq\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,z,i+1)\quad\forall x,y,z\in\chi$
-Next we show that these two properties can lead to a non-decreasing margin in the binary classification task, which suggests better performance.
+Next, we show that these two properties can lead to a monotonically decreasing upper bound on the test error in the binary classification task, which suggests improved performance.
 Consider a binary graph classification task and assume that the graph representations obtained at any iteration have a uniform norm.
@@
 Theorem 3.4
-Let $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ be a normalized iterative graph kernel that is monotonically decreasing and preserves order consistency.
-In the binary graph classification task with uniform graph representations, the margin between two classes in the representation space is non-decreasing w.r.t the iteration $i$, where the margin at iteration $i$ is defined as the shortest distance between two graph representations from different classes: margin=$\lVert\psi^{i}(x)-\psi^{i}(y)\rVert$, $\mathcal{Y}(x)\neq\mathcal{Y}(y)$.
+Let $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ be a normalized iterative graph kernel.
+Its generalization ability on the test set is provably strong when it decreases monotonically and preserves order consistency.
 Proof sketch.
-Suppose that at iteration $i$, the margin is defined by a pair of graphs, $x_{1}$ from class 1 and $y_{1}$ from class 2.
-In the next iteration, $i+1$, the margin is determined by another pair, $x_{2}$ and $y_{2}$.
-Two possibilities arise: (1) $x_{1}=x_{2}$ and $y_{1}=y_{2}$, or (2) $x_{1}\neq x_{2}$ or $y_{1}\neq y_{2}$.
+We consider a binary classification task where supervision is provided through labels and similarity orders.
+We demonstrate that the presence of these two properties enhances the generalizability of the learned iterative kernel.
+The proof is structured in two steps:
   1. (1)
-    If the pair defining the margin remains unchanged, the margin at iteration $i+1$ can only increase or stay the same, as the kernel is a monotonically decreasing function, indicating a reduction in similarity.
+    We first show that iterative kernels can learn the correct ranking given sufficient data during training, using the PAC-learning framework.
   2. (2)
-    We prove case 2 by contradiction, assuming the margin decreases at iteration $i+1$.
-This would imply that the kernel value for $x_{2}$ and $y_{2}$ increases at iteration $i+1$, which contradicts the fact that the kernel function is monotonically decreasing.
-Thus, in both cases, the margin does not decrease as the iterations progress.
-We provide the detailed proof in Appendix A.1
+    Next, we demonstrate that, once the correct ranking is learned, an iterative kernel learned with the properties of monotonic decrease and order consistency guarantees that the error rate of the graphs in the test set decreases as the number of iterations increases.
+Details of the proof can be found in Appendix A.
 3.3 Theoretical Verification with WL-based Kernels
@@
 $\tilde{\mathbb{K}}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime}\right)\geq\tilde{\mathbb{K}}_{WLOA}^{(h+1)}\left(\mathcal{G},\mathcal{G}^{\prime\prime}\right)$
-We include the complete proof in Appendix A.2 and A.3.
+We include the complete proof in Appendix B.1 and B.2.
 These findings imply that the WLOA kernel better preserves relational structures compared to the WL-subtree kernel, leading to improved classification performance, as supported by the literature (Kriege et al., 2016).
@@
 Thus, we aim to explore how to effectively preserve this principle within the GNN architectures.
-Figure 2: Computation of loss.
+Figure 2: Computation of Consistency loss.
 At each layer, pairwise distance matrix $\mathbf{D}$ is calculated using the normalized representations of graphs in a batch.
@@
 For ogbg-molhiv, we follow the official evaluator and use ROC-AUC as the evaluation metric.
-Detailed information about these datasets can be found in Appendix B.
+Detailed information about these datasets can be found in Appendix C.
 Model
@@
 To ensure a fair comparison, we maintain the same number of layers and layer sizes for both the base models and the models with our proposed consistency loss, ensuring the sharing of the same network architecture.
-Detailed information about hyperparameter tuning is provided in Appendix C.
+Detailed information about hyperparameter tuning is provided in Appendix D.
 5.2 Effectiveness of Consistency Loss
@@
 In addition, our method demonstrates satisfactory improvements on datasets with numerous classes (e.g., COIL-RAG) and large-scale datasets (e.g., ogbg-molhiv and Reddit-T), indicating that our approach is both flexible and scalable for handling complex and extensive datasets.
-Table 1: Classification performance on the TU and OGB datasets, with and without the consistency loss.
+Table 1: Classification performance on the TU, OGB and Reddit-T datasets, with and without the consistency loss.
 Highlighted cells indicate instances where the base GNN with the consistency loss outperforms the base GNN alone.
@@
 #nodes 29.87 29.68 39.06 284.32 19.77 13.00 74.49 3.01 25.50 23.93 GCN 73.96$\pm$ 2.37 74.04$\pm$ 3.09 73.24$\pm$ 6.93 74.92$\pm$ 2.66 75.40$\pm$ 2.97 55.07$\pm$1.24 81.72$\pm$0.84 91.72$\pm$1.65 72.86$\pm$1.90 76.00$\pm$0.44 +$\mathcal{L}_{\text{consistency}}$ 75.12$\pm$ 1.19 73.25$\pm$ 1.25 75.07$\pm$ 5.05 78.56$\pm$ 3.32 75.85$\pm$ 1.82 56.27$\pm$1.00 83.44$\pm$0.45 93.38$\pm$1.64 73.75$\pm$0.89 77.12$\pm$0.12 GIN 78.13$\pm$ 2.11 76.75$\pm$ 2.91 72.97$\pm$ 4.59 71.10$\pm$ 4.63 70.80$\pm$ 4.07 52.13$\pm$1.42 79.84$\pm$1.05 93.33$\pm$1.48 71.60 $\pm$2.36 77.50$\pm$0.16 +$\mathcal{L}_{\text{consistency}}$ 79.45$\pm$ 1.09 77.46$\pm$ 1.96 74.98$\pm$ 4.57 75.51$\pm$ 2.63 74.50$\pm$ 3.06 53.46$\pm$2.44 84.16$\pm$0.81 94.03$\pm$1.33 74.57$\pm$1.61 77.64$\pm$0.05 GraphSAGE 74.40$\pm$ 1.83 73.17$\pm$ 0.47 74.96$\pm$ 3.14 76.44$\pm$ 4.16 73.90$\pm$ 2.17 51.33$\pm$2.95 78.92$\pm$1.20 89.56$\pm$2.37 77.03$\pm$1.65 76.67$\pm$0.11 +$\mathcal{L}_{\text{consistency}}$ 78.26$\pm$ 1.08 74.10$\pm$ 2.10 76.40$\pm$ 3.12 77.50$\pm$ 3.38 74.75$\pm$ 3.06 54.27$\pm$1.24 82.12$\pm$0.78 92.31$\pm$1.32 78.60$\pm$1.44 77.57$\pm$0.05 GTransformer 75.72$\pm$2.69 74.79$\pm$1.82 73.33$\pm$4.80 75.42$\pm$3.22 72.20$\pm$3.49 53.33$\pm$1.12 80.36$\pm$0.56 83.74$\pm$3.17 76.81$\pm$1.34 76.75$\pm$0.12 +$\mathcal{L}_{\text{consistency}}$ 76.83$\pm$1.36 75.82$\pm$1.53 77.03$\pm$3.79 76.57$\pm$2.54 73.75$\pm$2.56 56.53$\pm$1.54 80.48$\pm$0.47 91.67$\pm$1.88 76.90$\pm$3.25 77.14$\pm$0.06 GMT 75.04$\pm$1.43 73.90$\pm$2.29 72.70$\pm$4.21 72.80$\pm$2.19 79.80$\pm$1.08 54.13$\pm$2.90 80.36$\pm$1.15 90.85$\pm$1.91 74.86$\pm$2.26 72.06$\pm$10.15 +$\mathcal{L}_{\text{consistency}}$ 75.52$\pm$1.07 75.20$\pm$0.95 74.86$\pm$2.03 73.14$\pm$2.28 79.60$\pm$1.91 54.80$\pm$1.42 82.80$\pm$0.61 92.00$\pm$1.43 76.00$\pm$1.99 77.19$\pm$0.14
-Furthermore, we analyze the complexity and scalability of our method on the TU Dataset, with additional details provided in Appendices D.1 and D.2.
-Additionally, we demonstrate the method’s potential to enhance performance significantly, even with a marginal increase in computational cost, as illustrated in Appendix E.
+Furthermore, we analyze the complexity and scalability of our method on the TU Dataset, with additional details provided in Appendices E.1 and E.2.
+Additionally, we demonstrate the method’s potential to enhance performance significantly, even with a marginal increase in computational cost, as illustrated in Appendix F.
 5.3 Effect of the Consistency Loss on Rank Correlation
 Table 2: Spearman correlation was computed for graph representations from consecutive layers on the TU datasets, both with and without consistency loss.
-Values with higher rank correlation are highlighted in grey.
+Values with higher rank correlation are highlighted.
 The consistency loss can enhance the rank correlation of graph similarities.
@@
 Specifically, we randomly sampled between 2 and 4 classes to construct new datasets from the original dataset and conducted classification tasks using both GCN and GCN with $\mathcal{L}_{\text{consistency}}$ on these newly constructed datasets.
-We report the mean test accuracy over five experiments for each subset, as presented in Table.3.
+We report the mean test accuracy over five experiments for each subset, as presented in Table 3.
 Table 3: Performance comparison across different subsets and the full set.
@@
 The above results show that the GCN model, enhanced with the $\mathcal{L}_{\text{consistency}}$ loss function, consistently outperforms the original version across different structural complexity groups, demonstrating the robustness and effectiveness of the proposed method.
-Furthermore, we evaluate the method’s efficiency under various complexity scenarios, as detailed in the Appendix D.3.
+Furthermore, we evaluate the method’s efficiency under various complexity scenarios, as detailed in the Appendix E.3.
 6 Related Work
@@
 Experiments show that our proposed method universally enhances the performance of backbone networks on graph classification benchmarks, providing new insights into bridging the gap between traditional kernel methods and GNN models.
-Appendix A Theoretical Verification with WL-based Kernels
-A.1 Non-decreasing Margin
-Suppose that at iteration $i$, graph $x_{1}$ in class 1 and graph $y_{1}$ in class 2 decide the margin and at iteration $i+1$, graph $x_{2}$ in class 1 and graph $y_{2}$ in class 2 decide the margin.
-There are two scenarios: (1) $x_{1}=x_{2}$ and $y_{1}=y_{2}$, (2) $x_{1}\neq x_{2}$ or $y_{1}\neq y_{2}$. For case 1, margin at iteration $i$ is given by:
-$\begin{split}\lVert\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1}))\rVert&=\sqrt{(\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1})))^{T}(\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1})))}\\ &=\sqrt{2-2\langle\phi(\psi^{i}(x_{1})),\phi(\psi^{i}(y_{1}))\rangle}=\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{1},y_{1},i)}.\end{split}$ (1)
-Since $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ is monotonically decreasing, we have:
-$\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{1},y_{1},i)}\leq\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{1},y_{1},i+1)}=\lVert\phi(\psi^{i+1}(x_{1}))-\phi(\psi^{i+1}(y_{1}))\rVert.$ (2)
-This suggests that the margin does not decrease.
-For case 2, we prove by contradiction and first assume that the margin decreases. Since the margin decreases, we have:
-$\lVert\phi(\psi^{i+1}(x_{2}))-\phi(\psi^{i+1}(y_{2}))\rVert<\lVert\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1}))\rVert.$ (3)
-At iteration $i$, since $x_{1}$ and $y_{1}$ decide the margin, we have:
-$\lVert\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1}))\rVert\leq\lVert\phi(\psi^{i}(x_{2}))-\phi(\psi^{i}(y_{2}))\rVert.$ (4)
-Combining Equation 3 and Equation 4, we have:
-$\lVert\phi(\psi^{i+1}(x_{2}))-\phi(\psi^{i+1}(y_{2}))\rVert<\lVert\phi(\psi^{i}(x_{2}))-\phi(\psi^{i}(y_{2}))\rVert.$ (5)
-This is equivalent to:
-$\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{2},y_{2},i+1)}<\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{2},y_{2},i)},$ (6)
+Appendix A Proof of Theorem 3.4
+A.1 Learning Correct Similarity Order with Iterative Kernels
+In this section, we first prove that the iterative kernel with bounded feature mapping is capable of learning the pairwise similarity order given sufficient training data, as shown by the PAC-learning framework.
+We consider a dataset consisting of instances $x\in\chi$, where $\chi$ is the input space.
+The dataset is denoted as $\chi=\left\{x_{1},x_{2},\ldots,x_{n}\right\}$.
+We employ a kernel function $\mathbb{K}:\chi\times\chi\rightarrow\mathbb{R}$ with an associated feature mapping $\phi:\chi\rightarrow\mathcal{H}_{\mathbb{K}}$ such that $\mathbb{K}(x,y)=\langle\phi(x),\phi(y)\rangle_{\mathcal{H}_{\mathbb{K}}}$.
+Here, $\mathcal{H}_{\mathbb{K}}$ is the Reproducing Kernel Hilbert Space (RKHS) associated with $\mathbb{K}$, and $\langle\cdot,\cdot\rangle_{\mathcal{H}{\mathbb{K}}}$ denotes the inner product in $\mathcal{H}_{\mathbb{K}}$.
+We assume that the feature mappings are bounded, i.e., $\|\phi(x)\|_{\mathcal{H}_{\mathbb{K}}}\leq R,\forall x\in\chi$ for some $R>0$.
+Using this setup, we focus on an arbitrary fixed data point $x_{i}$ and compute its similarity with other data points $x_{j}(j\neq i)$ using the kernel function:
+$S=\mathbb{K}\left(x_{i},x_{j}\right)=\left\langle\phi\left(x_{i}\right),\phi\left(x_{j}\right)\right\rangle_{\mathcal{H}_{\mathbb{K}}}$ (1)
+Thus, we define the function $f$ as $f\left(x_{j}\right)=\left\langle w,\phi\left(x_{j}\right)\right\rangle_{\mathcal{H}_{\mathbb{K}}}$ where $w=\phi\left(x_{i}\right)$.
+Since $w=\phi\left(x_{i}\right)$ and $\left\|\phi\left(x_{i}\right)\right\|_{\mathcal{H}_{\mathbb{K}}}\leq R$, it follows that $\|w\|_{\mathcal{H}_{\mathbb{K}}}\leq R$.
+So the function $f$ representing the similarity between $x_{i}$ and any other point $x_{j}$ can be formulated as:
+$f\left(x_{j}\right)=\left\langle\phi\left(x_{i}\right),\phi\left(x_{j}\right)\right\rangle_{\mathcal{H}_{\mathbb{K}}}=\mathbb{K}\left(x_{i},x_{j}\right)=S_{ij}.$ (2)
+Since we are dealing with a ranking problem, we use the hinge loss function for pairs $\left(x_{j},x_{k}\right)(j,k\neq i)$ :
+$\ell_{\text{hinge }}\left(f,x_{j},x_{k}\right)=\max\left(0,1-s_{jk}^{*}\left(f\left(x_{j}\right)-f\left(x_{k}\right)\right)\right),$ (3)
+where $s_{jk}^{*}=\operatorname{sign}\left(f^{*}\left(x_{j}\right)-f^{*}\left(x_{k}\right)\right)$, and $f^{*}$ represents the ground-truth ranking function.
+Thus, the expected loss over all pairs is given by:
+$L(f)=\mathbb{E}_{\left(x_{j},x_{k}\right)}\left[\ell_{\text{hinge }}\left(f,x_{j},x_{k}\right)\right],$ (4)
+and the empirical loss over the sample $S$ is given by:
+$\hat{L}(f)=\frac{2}{(n-1)(n-2)}\sum_{\begin{subarray}{c}j<k\\ j,k\neq i\end{subarray}}\ell_{\text{hinge }}\left(f,x_{j},x_{k}\right),$ (5)
+since there are $\frac{(n-1)(n-2)}{2}$ pairs among the $n-1$ data points excluding $x_{i}$.
+To assess how well the learned function $f$ generalizes to test data, we derive a generalization bound using Rademacher complexity.
+We define the function class for pairs by:
+$\mathcal{F}_{\text{pair }}=\left\{f\left(x_{j},x_{k}\right)=f\left(x_{j}\right)-f\left(x_{k}\right)\right\}.$ (6)
+which represents the similarity difference between $x_{j}$ and $x_{k}$ given the reference graph $x_{i}$.
+Given that $f(x)=\langle w,\phi(x)\rangle$ with $\|w\|\leq R$ and $\|\phi(x)\|\leq R$, we can bound $f\left(x_{j},x_{k}\right)$ by:
+$\left|f\left(x_{j},x_{k}\right)\right|=\left|f\left(x_{j}\right)-f\left(x_{k}\right)\right|\leq\left|f\left(x_{j}\right)\right|+\left|f\left(x_{k}\right)\right|\leq 2R^{2},$ (7)
+since $|f(x)|=|\langle w,\phi(x)\rangle|\leq\|w\|\cdot\|\phi(x)\|\leq R\cdot R=R^{2}$.
+Given $f\left(x_{j},x_{k}\right)$ takes values in the range $\left[-2R^{2},2R^{2}\right]$.
+Therefore, following the PAC-learning framework in [Mohri, 2018], we derive the following generalization bound.
+For any $\delta>0$, with probability at least $1-\delta$ over a sample $S$ of size $m$ drawn, the following holds for all $f\in$ $\mathcal{F}_{\text{pair }}:$
+$L(f)\leq\hat{L}(f)+2\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)+M\sqrt{\frac{\ln(1/\delta)}{2m}},$ (8)
+where:
+  • ^$\hat{\mathfrak{R}}_{m}\left(\mathcal{F}_{\text{pair }}\right)$ is the empirical Rademacher complexity of $\mathcal{F}_{\text{pair }}$.
+  • $M=1+2R^{2}$ is an upper bound on the hinge loss function $\ell_{\text{hinge }}$.
+  • $m=\frac{(n-1)(n-2)}{2}$ is the number of pairs.
+  • $\delta$ is the confidence level.
+It is observed that as $m$ increases, the term $\sqrt{\frac{\ln(1/\delta)}{2m}}$ decreases at a rate proportional to $\sqrt{1/m}$.
+Next, we proceed by proving that the Rademacher complexity is bounded and decreases as $m$ increases.
+Theorem A.1
+Assuming $\|\phi(x)\|_{\mathcal{H}_{K}}\leq R$ for all $x\in\chi$, the empirical Rademacher complexity of the function class $\mathcal{F}_{\text{pair}}$ can be bounded as:
+^$\hat{\mathfrak{R}}_{S}(\mathcal{F}_{\text{pair}})\leq\frac{2R^{2}}{\sqrt{m}}.$ (9)
+Proof 1
+Based on the definition of Rademacher complexity, for function class $\mathcal{F}_{\text{pair }}$ with a sample size $m$, the complexity ^$\hat{\mathfrak{R}}_{S}(\mathcal{F}_{\text{pair}})$ is defined as:
+^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)=\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}_{\text{pair }}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}f\left(x_{j_{i}},x_{k_{i}}\right)\right],$ (10)
+where $\sigma=\left(\sigma_{1},\ldots,\sigma_{m}\right)$ are independent Rademacher variables (i.e., each $\sigma_{k}$ takes the value +1 or -1 with equal probability), and $\left\{\left(x_{j_{i}},x_{k_{i}}\right)\right\}_{i=1}^{m}$ are pairs sampled from the data.
+We can express ^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)$ as:
+^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)=\frac{1}{m}\mathbb{E}_{\sigma}\left[\sup_{\|w\|\leq R}\left\langle w,\sum_{i=1}^{m}\sigma_{i}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\rangle\right].$ (11)
+Applying the Cauchy-Schwarz inequality:
+$\langle w,v\rangle_{\mathcal{H}_{\mathbb{K}}}\leq\|w\|_{\mathcal{H}_{\mathbb{K}}}\|v\|_{\mathcal{H}_{\mathbb{K}}}\leq R\|v\|_{\mathcal{H}_{\mathbb{K}}}$ (12)
+we have:
+^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)\leq\frac{1}{m}R\cdot\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\|\right].$ (13)
+Using Jensen’s inequality in the form $\mathbb{E}[\|X\|]\leq\left(\mathbb{E}\left[\|X\|^{2}\right]\right)^{1/2}$, we obtain:
+^$\hat{\mathfrak{R}}_{S}(\mathcal{F}_{\text{pair }})\leq\frac{R}{m}\left(\mathbb{E}_{\sigma}\left[\left\|\sum_{k=1}^{m}\sigma_{k}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\|_{\mathcal{H}_{K}}^{2}\right]\right)^{1/2}$ (14)
+Since $\sigma_{i}$ are independent Rademacher variables, we have $\mathbb{E}\sigma\left[\sigma_{i}^{2}\right]=1$. Therefore:
+$\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\|^{2}\right]$ $\leq\sum_{i=1}^{m}\mathbb{E}_{\sigma}\left[\sigma_{i}^{2}\right]\left\|\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right\|^{2}$ (15)
+$=\sum_{i=1}^{m}\left\|\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right\|^{2}.$
+Given that $\|\phi(y)\|_{\mathcal{H}_{\mathbb{K}}}\leq R$ , we have:
+$\left\|\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right\|^{2}\leq(2R)^{2}=4R^{2}.$ (16)
+Therefore,
+$\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\|\right]\leq\sqrt{m\cdot 4R^{2}}=2R\sqrt{m}.$ (17)
+Thus
+^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)\leq\frac{1}{m}R\cdot 2R\sqrt{m}=\frac{2R^{2}}{m}\sqrt{m}=\frac{2R^{2}}{\sqrt{m}}.$ (18)
+Therefore, the empirical Rademacher complexity decreases with $m$.
 […]
-A.2 Proof of Theorem 3.5
+Thus, consider Equation 8:
+  • The term $\sqrt{\frac{\ln(1/\delta)}{2m}}$ decreases proportionally to $1/\sqrt{m}$.
+  • Empirical Rademacher complexity ^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)\leq\frac{2R^{2}}{\sqrt{m}}$ also decreases proportionally to $1/\sqrt{m}$.
+Therefore, the generalization gap between $L(f)$ and $\hat{L}(f)$ decreases as $m$ increases, demonstrating that the kernel-based ranking function is PAC-learnable.
+A.2 Linking Two Key Properties to Generalizability in Binary Classification
+Building on the setup and conclusions from the previous section, we establish that the kernel can learn the correct ranking with sufficient training data.
+Here, we investigate how, once the correct ranking is achieved, an iterative kernel with monotonic decrease and order consistency properties ensures a decreasing error rate for test set graphs as iterations progress in binary classification tasks.
+To formalize this, we define a probability space over the whole graph space where addition and scalar multiplication are defined.
+Let $\Omega$ be the set of all possible data points (graphs).
+Each data point $g_{i}\in\Omega$ represents a possible outcome, and the event of observing $g_{i}$ is the singleton set $\left\{g_{j}\right\}\subseteq\Omega$.
+By considering all subsets of $\Omega$, we construct an event space $\mathcal{F}$.
+Assuming the existence of an underlying probability distribution over these outcomes, we obtain the probability space $(\Omega,\mathcal{F},P)$.
+Using this framework, let $\mu_{x}$ and $\mu_{y}$ represent the unknown mean graph in class 1 and class 2, respectively.
+Let $\mathbb{K}_{gt}(\cdot,\cdot)\in[0,1]$ represent the ground-truth kernel function for pairs of graphs.
+Without loss of generality, we analyze the error rate for graphs belonging to class 1; the error rate for class 2 can be derived in a similar manner.
+We define two random variables associated with the ground-truth kernel: $\mathbf{s}_{x}=\mathbb{K}_{gt}({\mu_{x},x)}$ and $\mathbf{s}_{y}=\mathbb{K}_{gt}({\mu_{y},x)}$, where $x$ is a randomly sampled instance from class 1.
+The terms $\mathbf{s}_{x}(x)$ and $\mathbf{s}_{y}(x)$ denote specific realizations of $\mathbf{s}_{x}$ and $\mathbf{s}_{y}$, respectively, for a given realization of the graph $x$ belonging to class 1.
+Since $\mathbf{s}_{x}=0$ implies that there exists graph from class 1 that is completely different from $\mu_{x}$, which is counterintuitive, we assume that the probability density function $p(u)$, $u\in[0,1]$, of $\mathbf{s}_{x}$ has the following asymptotic property on the right neighborhood of $u=0$:
+$\lim_{u\to 0^{+}}\frac{p(u)}{u}=C,$ (19)
+where constant $C$ satisfies that $0\leq C<+\infty$.
+Let $\mathbf{\bar{s}}^{(i)}_{x}(x)=\mathbb{K}^{(i)}(\bar{\mu}_{x}^{(i)},x)$ and $\mathbf{\bar{s}}^{(i)}_{y}(x)=\mathbb{K}^{(i)}(\bar{\mu}_{y}^{(i)},x)$ represent the estimated realizations of $\mathbf{s}_{x}$ and $\mathbf{s}_{y}$ at iteration $i$, respectively, for a given graph realization $x$.
+$\bar{\mu}_{x}^{(i)}$ and $\bar{\mu}_{y}^{(i)}$ are estimates for $\mu_{x}$ and $\mu_{y}$, respectively.
+We assume that $\bar{\mu}_{x}^{(i)}$ and $\bar{\mu}_{y}^{(i)}$ are obtained through operations on the training graphs, where these operations are closed in the graph space.
+Due to the capability of the iterative kernels, we further assume that as the iteration $i\to\infty$, the sequences $\{\bar{\mu}_{x}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$, $\{\bar{\mu}_{y}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$ and $\{\mathbb{K}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$ converge to $\mu_{x}$, $\mu_{y}$ and $\mathbb{K}_{gt}$, respectively.
+We assume that the convergence rate of $\{\mathbb{K}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$ is much slower than the convergence rate of $\{\bar{\mu}_{x}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$ and $\{\bar{\mu}_{y}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$.
+Furthermore, every function in the series of function $\{\mathbb{K}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$ is continuous.
+Denote the training data for class 1 and class 2 as $\left\{x_{1},x_{2},\ldots,x_{n}\right\}\subset\Omega$ and $\left\{y_{1},y_{2},\ldots,y_{m}\right\}\subset\Omega$, respectively.
+Since the learned kernels $\mathbb{K}^{(i)}$ preserve the similarity order, we can assume, without loss of generality, the following empirical ordering of the training graphs for the kernel at the $i$-th iteration:
+$\mathbf{\bar{s}}^{(i)}_{x}({x_{1})}<\mathbf{\bar{s}}^{(i)}_{x}({x_{2})}<\cdots<\mathbf{\bar{s}}^{(i)}_{x}({x_{n})}$ (20)
+In this setting, for an arbitrary test graph $g$ sampled from the graph space, the decision boundary of the kernel learned during the $i$-th iteration is defined as follows:
+$\forall g:\,\text{if}\;\mathbf{\bar{s}}^{(i)}_{x}(g)>\mathbf{\bar{s}}^{(i)}_{y}(g)\;\text{then}\,g\,\text{is classified as class 1}$ (21)
+$\forall g:\,\text{if}\;\mathbf{\bar{s}}^{(i)}_{x}(g)<\mathbf{\bar{s}}^{(i)}_{y}(g)\;\text{then}\,g\,\text{is classified as class 2}$
+We assume that $\mathbf{\bar{s}}^{(i)}_{x}(g)$ and $\mathbf{\bar{s}}^{(i)}_{y}(g)$, as well as $\mathbf{{s}}_{x}(g)$ and $\mathbf{{s}}_{y}(g)$, are not equal for any of the training data points.
+Next, we analyze the misclassification rate for graphs in class 1.
+Let the integer $k$ satisfies: $\mathbf{\bar{s}}^{(i)}_{x}(x_{k})<\mathbf{\bar{s}}^{(i)}_{y}(x_{k})$, while $\mathbf{\bar{s}}^{(i)}_{x}(x_{k+1})>\mathbf{\bar{s}}^{(i)}_{y}({x_{k+1})}$ or a given iteration $i$, which means $x_{k}$ is the misclassified training graph in class 1 closest to the decision boundary.
+If the training error is 0, we take $x_{0}=y_{1}$.
+We note that the integer $k$ is independent of the iteration $i$ due to the consistency principle.
+Next, we demonstrate this using the following lemma.
+Lemma A.2
+For any sufficiently large iteration number $i$, there exists a fixed index $k$ such that $\mathbf{\bar{s}}^{(i)}_{x}(x_{k})<\mathbf{\bar{s}}^{(i)}_{y}(x_{k})$, while $\mathbf{\bar{s}}^{(i)}_{x}(x_{k+1})>\mathbf{\bar{s}}^{(i)}_{y}({x_{k+1})}$.
+Proof 2
+Since the series $\{\bar{\mu}_{x}^{(i)}\}_{i=1}^{+\infty}$ converges to $\mu_{x}$, it satisfies Cauchy’s convergence criterion.
+Specifically, $\forall\delta>0$, $\exists N_{1}>0$, such that when $i>N_{1}$,
+$|\bar{\mu}_{x}^{(i+1)}-\bar{\mu}_{x}^{(i)}|<\delta.$ (22)
+Since we have assumed that $\mathbb{K}^{(i)}(\cdot,\cdot)$ are continuous: $\forall\epsilon>0$, $\exists\delta>0$, $\exists N_{2}>0$, such that for any $x$ in class 1, $\forall i>N_{2}$ and $|\bar{\mu}_{x}^{(i+1)}-\bar{\mu}_{x}^{(i)}|<\delta$, we have:
+$\lvert\mathbb{K}^{(i+1)}(\bar{\mu}_{x}^{(i+1)},x)-\mathbb{K}^{(i+1)}(\bar{\mu}_{x}^{(i)},x)\rvert<\epsilon.$ (23)
+Based on Equation 23, we have:
+[…]
+Since the consistency principle holds, $\mathbb{K}^{(i+1)}(\bar{\mu}_{x}^{(i)},x_{k})-\mathbb{K}^{(i+1)}(\bar{\mu}_{y}^{(i)},x_{k})$ and $\mathbf{\bar{s}}^{(i)}_{x}(x_{k})-\mathbf{\bar{s}}^{(i)}_{y}(x_{k})$ have the same sign.
+If $\mathbf{\bar{s}}^{(i)}_{x}(x_{k})-\mathbf{\bar{s}}^{(i)}_{y}(x_{k})<0$, then $\mathbb{K}^{(i+1)}(\bar{\mu}_{x}^{(i)},x_{k})-\mathbb{K}^{(i+1)}(\bar{\mu}_{y}^{(i)},x_{k})<0$.
+For a sufficiently large $N$, we select an $\epsilon>0$ such that $\epsilon<\min_{\{i>N\}}\frac{\lvert\mathbb{K}^{(i+1)}(\bar{\mu}_{x}^{(i)},x_{k})-\mathbb{K}^{(i+1)}(\bar{\mu}_{y}^{(i)},x_{k})\rvert}{2}$.
+Given our assumption that no training data point $x_{k}$ lies precisely on the decision boundary, such an $\epsilon$ is guaranteed to exist.
+Consequently, Equation 24 simplifies to:
+$\mathbf{\bar{s}}^{(i+1)}_{x}(x_{k})-\mathbf{\bar{s}}^{(i+1)}_{y}(x_{k})<2\epsilon+\mathbb{K}^{(i+1)}(\bar{\mu}_{x}^{(i)},x_{k})-\mathbb{K}^{(i+1)}(\bar{\mu}_{y}^{(i)},x_{k})<0.$ (25)
+Thus, $\mathbf{\bar{s}}^{(i+1)}_{x}(x_{k})-\mathbf{\bar{s}}^{(i+1)}_{y}(x_{k})$ have the same sign as $\mathbf{\bar{s}}^{(i)}_{x}(x_{k})-\mathbf{\bar{s}}^{(i)}_{y}(x_{k})$.
+If $\mathbf{\bar{s}}^{(i)}_{x}(x_{k})-\mathbf{\bar{s}}^{(i)}_{y}(x_{k})>0$, the same conclusion can be derived in a similar manner.
+Since the number of training samples is finite, we can find a sufficiently large iteration such that Equation 25 holds for all $x_{k}$ after that iteration.
+Thus, some point between $\mathbf{\bar{s}}^{(i)}_{x}(x_{k})$ and $\mathbf{\bar{s}}^{(i)}_{x}(x_{k+1})$ is the decision boundary learned at the $i$-th iteration.
+Based on our assumption that the convergence rate of $\{\mathbb{K}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$ is much slower than the convergence rate of $\{\bar{\mu}_{x}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$ and $\{\bar{\mu}_{y}^{(i)}(\cdot,\cdot)\}_{i=1}^{+\infty}$, $\mathbf{\bar{s}}^{(i)}_{x}(x_{k})$ is asymptotically monotonically decreasing.
+$\mathbf{\bar{s}}^{(i)}_{x}(x_{k})>\mathbf{\bar{s}}^{(i+1)}_{x}(x_{k}),\text{when }i\to+\infty$ (26)
+Thus,
+$\mathbf{\bar{s}}^{(i)}_{x}(x_{k})\to\mathbb{K}_{gt}\left(\mu_{x},x_{k}\right)^{+}\ \text{when }i\to+\infty\implies\mathbf{\bar{s}}^{(i)}_{x}(x_{k})>\mathbb{K}_{gt}\left(\mu_{x},x_{k}\right)$ (27)
+Thus, $\forall x\in$ class 1, the probability of misclassifying $x$ is given by: $\mathbb{P}\{\mathbf{\bar{s}}^{(i)}_{x}(x_{k})<\mathbb{K}_{gt}(\mu_{x},x_{k})\}$. Specifically, we have:
+$\mathbb{P}\{\mathbf{s}_{x}<\mathbb{K}_{gt}(\mu_{x},x_{k})\}\leq\mathbb{P}\{\mathbf{s}_{x}\leq\mathbf{\bar{s}}^{(i)}_{x}(x_{k})\}$ (28)
+In the following theorem, we show that the probability of misclassifying any graph from class 1 can be bounded using Markov’s inequality.
+Theorem A.3
+If the iterative kernel decreases monotonically and preserves the similarity order, the probability of misclassifying a graph $x\in\chi$ from class 1 has an upper bound that decreases monotonically as the number of iterations increases.
+Proof 3
+To upper bound the error rate, we begin by applying the transformation $\frac{1}{\mathbf{s}_{x}}-1$ to the original random variable $\mathbf{s}_{x}$. The expectation of this transformed variable is:
+[…]
+[…]
+Since $u=0$ is singular point, we reformulate the improper integral as:
+[…]
+Here, since $\lim_{u\to 0^{+}}\frac{p(u)}{u}=C$, the limit exists and thus the given integral converges, ensuring the existence of the expectation.
+With this result, we can apply Markov’s inequality to the transformed random variable $\frac{1}{\mathbf{s}_{x}}-1$.
+For any nonnegative $a$, we have:
+$\mathbb{P}\left[(\frac{1}{\mathbf{s}_{x}}-1)\geqslant a\right]\leqslant\frac{\mathbb{E}\left(\frac{1}{\mathbf{s}_{x}}-1\right)}{a}$ (30)
+Thus
+$\mathbb{P}\left[\mathbf{s}_{x}\leqslant\frac{1}{1+a}\right]\leqslant\frac{\mathbb{E}\left(\frac{1}{\mathbf{s}_{x}}-1\right)}{a}$ (31)
+Let $a=\frac{1}{\mathbf{\bar{s}^{(i)}}_{x}(x_{k})}-1$. Referring to Equation 28, we have:
+$\mathbb{P}\{\mathbf{s}_{x}<\mathbb{K}_{gt}(\mu_{x},x_{k})\}\leqslant\mathbb{P}\{\mathbf{s}_{x}\leq\mathbf{\bar{s}}^{(i)}_{x}(x_{k})\}\leqslant\frac{\mathbb{E}\left(\frac{1}{\mathbf{s}_{x}}-1\right)}{\frac{1}{\mathbf{\bar{s}^{(i)}}_{x}(x_{k})}-1}$ (32)
+Since $\mathbf{\bar{s}^{(i)}}_{x}(x_{k})$ decreases monotonically and $\mathbb{E}\left(\frac{1}{\mathbf{s}_{x}}-1\right)$ is a constant, it follows that $\mathbb{P}\{\mathbf{s}_{x}<\mathbb{K}_{gt}(\mu_{x},x_{k})\}$ has a monotonically decreasing upper bound.
+Appendix B Theoretical Verification with WL-based Kernels
+B.1 Proof of Theorem 3.5
 Following Equation 2.4, the normalized WL-subtree kernel after $h$ iterations can be expressed as:
@@
 […]
-A.3 Proof of Theorem 3.6
-A.3.1 Monotonic Decrease in WLOA Kernel
+B.2 Proof of Theorem 3.6
+B.2.1 Monotonic Decrease in WLOA Kernel
 The WLOA kernel after $h$ iterations can be computed as follows:
@@
 Therefore, we get $D\geq 0$. The value of Equation 33 is non-negative, indicating that WLOA is monotonically decreasing.
-A.3.2 Order Consistency in WLOA Kernel
+B.2.2 Order Consistency in WLOA Kernel
 Assume two graph pairs satisfy the following relation:
@@
 […]
-Appendix B Dataset
+Appendix C Dataset
 In this section, we provide the statistics of the datasets used, as shown in Table 5.
@@
 Reddit-T Reddit Thread Classification 2 203088 23.93 24.99 -
-Appendix C Detailed Set-Up
+Appendix D Detailed Set-Up
 TU Dataset
@@
 Each model is trained and evaluated over 5 independent runs, with the mean and standard deviation of the results recorded for performance comparison.
-Appendix D Complexity& Scalability
-D.1 Time Complexity
+Appendix E Complexity& Scalability
+E.1 Time Complexity
 We present the time complexity analysis for our proposed consistency loss.
@@
 GCN+$\mathcal{L}_{\text{consistency}}$ 8.861 6.358 5.529 5.382 5.252
-D.2 Space Complexity
+E.2 Space Complexity
 Next, we present the space complexity analysis for our consistency loss.
@@
 Cost Increase (%) 2.70 4.96 3.68 3.60 2.55
-D.3 Efficiency on Different Task and Structural Complexities
+E.3 Efficiency on Different Task and Structural Complexities
 Task Complexity
@@
 GCN+$\mathcal{L}_{\text{consistency}}$ $0.0371$ $0.0378$ $0.0392$
-Appendix E Efficient Consistent Learning
+Appendix F Efficient Consistent Learning
 To further minimize the overhead of our proposed consistency loss, we examined a scenario where the consistency loss, denoted as $\mathcal{L}_{FL}$, was only applied to the first and last layers.
@@
 This finding suggests that our proposed approach can be accelerated with minimal additional computational cost while still enhancing performance, thereby validating the effectiveness of the consistent learning principle.
-Appendix F Similarity/Difference with Contrastive learning
+Appendix G Similarity/Difference with Contrastive learning
 In this section, we discuss the similarities and differences between our method and graph contrastive learning.
@@
 This underscores the significant differences between our approach and regular GCL methods.
-Appendix G Boarder Impact
+Appendix H Boarder Impact
 This paper aims to advance the field of graph learning by proposing a model-agnostic consistency learning framework.
@@
 Additionally, we do not foresee any direct negative societal or ethical consequences stemming from our work.
-Appendix H Limitation
+Appendix I Limitation
 One limitation of our work is that the method involves additional computational costs, especially during large batch training processes.